### v1

# Tight Robustness Certificates and Wasserstein Distributional Attacks for Deep Neural Networks

Abstract
Wasserstein distributionally robust optimization (WDRO) provides a framework for adversarial robustness, yet existing methods based on global Lipschitz continuity or strong duality often yield loose upper bounds or require prohibitive computation. In this work, we address these limitations by introducing a primal approach and adopting a notion of exact Lipschitz certificate to tighten this upper bound of WDRO. In addition, we propose a novel Wasserstein distributional attack (WDA) that directly constructs a candidate for the worst-case distribution. Compared to existing point-wise attack and its variants, our WDA offers greater flexibility in the number and location of attack points. In particular, by leveraging the piecewise-affine structure of ReLU networks on their activation cells, our approach results in an exact tractable characterization of the corresponding WDRO problem. Extensive evaluations demonstrate that our method achieves competitive robust accuracy against state-of-the-art baselines while offering tighter certificates than existing methods. Our code is available at https://github.com/OLab-Repo/WDA
1 Introduction
Modern deep networks achieve remarkable accuracy yet remain fragile to distribution shift and adversarial perturbations (Szegedy et al., 2014; Goodfellow et al., 2014; Kurakin et al., 2018; Hendrycks & Dietterich, 2019; Ovadia et al., 2019; Taori et al., 2020; Koh et al., 2021), raising concerns about their reliability in deployment. A principled avenue for robustness is Wasserstein distributionally robust optimization (WDRO, Mohajerin Esfahani & Kuhn 2018; Gao & Kleywegt 2023), which controls worst-case test risk over an ambiguity set within a Wasserstein ball around the empirical distribution and admits tight dual characterizations from optimal transport (Villani, 2008; Santambrogio, 2015). While numerous defenses have been proposed, a fundamental gap persists between theoretical robustness certificates and practical adversarial evaluation: existing Lipschitz-based certificates often provide loose upper bounds that vastly overestimate the true worst-case loss (Virmaux & Scaman, 2018), while standard attacks restrict perturbations to fixed-radius balls around individual points (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018). This mismatch stems from two limitations: certificates typically rely on global worst-case analysis that ignores the actual network geometry traversed by data, and attacks consider only point-wise perturbations rather than distributional shifts permitted by Wasserstein threat models (Singh et al., 2018; Gao & Kleywegt, 2023). The discrepancy is particularly pronounced for modern architectures with ReLU activations, where the network behaves as a piecewise-affine function whose local properties vary dramatically across regions (Jordan & Dimakis, 2020), and those with smooth activations (GELU, SiLU/Swish) exhibit complex nonlinear geometry (Hendrycks & Gimpel, 2016; Ramachandran et al., 2017; Elfwing et al., 2018). In this work, we aim to address both sides of this gap: our contributions can be summarized as follows.
  1. 1.
    For a class of networks with Rectified Linear Unit (ReLU) activations (Nair & Hinton, 2010), we analyze the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem by connecting with the tight Lipschitz constant studied in Jordan & Dimakis, 2020. Our analysis is based on the classical underlying piecewise-affine structure of ReLU networks: on any strict ReLU cell, the logit map $\theta(\cdot)$ is affine with a constant input-logit Jacobian $J_{D}$. Our contribution is to leverage this structure for WDRO, which requires combining the Lipschitz constant of the logit map and the sensitivity of the softmax cross-entropy, or the DLR loss. Our first theoretical result yields an upper bound of WDRO induced by $\bm{L}\triangleq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}$, where $J_{\bm{D}}$ is general Jacobian of the logit map. (See 3.1 for precise definition of $J_{\bm{D}}$.) In addition, we derive a lower bound of WDRO by constructing a concrete and finite worst-case distribution. (See equation 16 for the explicit formulation.) This worst-case distribution is constructed by perturbing the empirical sample along the direction which the logit map is most varied. Moreover, we provide a sufficient condition where our lower and upper bounds match, and simulate an instance to illustrate this tightness, see Figure 2.
  2. 2.
    We further analyze the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem for a class of MLP with smooth activation and cross-entropy loss. Unlike ReLU activation or DLR loss, which might create degeneration edges, the chain rule is readily applied in this case, and the Lipschitz constant of the loss is naturally computed by estimating its gradients. Similar to the analysis of the ReLU networks, we obtain the upper bound of the WDRO as $\bm{L}\triangleq 2^{1/s}\max_{x\in\mathcal{X}}\left\|\nabla\theta(x)\right\|_{{r}\to{s}}$ while the worst-case distribution and lower bound are constructed similar to the ReLU networks.
  3. 3.
    Finally, we bridge the gap between WDRO theory and adversarial evaluation by introducing the Wasserstein Distributional Attack (WDA), which directly constructs adversarial distributions within the Wasserstein ball rather than restricting to point-wise perturbations. Unlike existing attacks that place all adversarial examples on the $\epsilon$-ball boundary, WDA flexibly interpolates between point-wise ($\kappa=1$) and truly distributional attacks ($\kappa>1$) by supporting adversarial distributions on $2N$ points. This offers a complementary perspective to strong baselines such as AutoAttack and the RobustBench leaderboard (Croce & Hein, 2020; Croce et al., 2021). Empirically, WDA with $\kappa=2$ consistently finds stronger adversarial examples than state-of-the-art methods across diverse settings: achieving lower robust accuracy than APGD-DLR on CIFAR-10/100 with WideResNet backbone (Zagoruyko & Komodakis, 2016) on both $\ell_{\infty}$ and $\ell_{2}$ perturbations. When integrated into the Adaptive Auto Attack framework, our method matches or exceeds the ensemble performance of A3. These results demonstrate that the distributional perspective not only provides tighter theoretical certificates but also yields more effective attacks, validating our claim that existing robustness evaluations underestimate vulnerability by restricting to $\Omega_{\infty}$ rather than the larger $\Omega_{1}$ ambiguity set assumed by certificates.
Figure 1: Left: Wasserstein ambiguity ball $\Omega_{p}=\left\{\mathbb{P}\colon\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\right\}$ inclusion and its admissible attacks. Our proposed Wasserstein Distributional Attack (WDA) with $\kappa\geq 1$ includes its special case $\kappa=1$ as a point-wise attack, and produces a distributional attack when $\kappa>1$. Note that most of the existing tight certificates estimated an upper bound of WDRO w.r.t. $\Omega_{p=1}$, not $\Omega_{p=\infty}$. Right: Visualization of point-wise attack (NN adversarial samples) versus our WDA ($2N$ adversarial samples). Our WDA allows not only a larger number of supports but also a wider range of perturbations.
2 Preliminaries
Notations
We denote basis vector as $\bm{e}_{k}$; indicator function as $\bm{1}_{\{\cdot\}}$; dirac measure as $\bm{\delta}_{z}$; input dimension $n$, and output dimension as $K$. An empirical dataset is denoted $\{Z^{(1)},\dots,Z^{(N)}\}$ with $Z=(x,y)\in\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{K}$; empirical distribution $\mathbb{P}_{N}=\sum_{i}\mu_{i}\bm{\delta}_{Z^{(i)}}$ with $Z^{(i)}=(x^{(i)},y^{(i)}=\bm{e}_{k_{i}})$. Norms $\|\cdot\|_{r}$ and $\|\cdot\|_{s}$ are dual with 1/r+1/s=11/{r}+1/{s}=1. For a matrix $A$, $\|A\|_{{r}\to{s}}=\sup_{\|u\|_{r}=1}\|Au\|_{s}$. Rectifier $\cdot]_{+}=\max{\{0,\cdot\}}$. Recession cone $\mathrm{rec}(\cdot)$. Interior set $\mathrm{int}(\cdot)$. Ground cost $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$. Cross-entropy loss $\ell(x,y;\theta)=-\sum_{k=1}^{K}y_{k}\log\operatorname{softmax}(\theta(x))_{k}$. Analogous DLR loss as defined in Croce & Hein (2020). We define the dual-norm maximizer $\mathcal{M}_{r}$ by
$\mathcal{M}_{r}\colon g\mapsto\arg\max_{h}\left\{\langle g,h\rangle\mid\|h\|_{r}=1\right\}=\left\{\begin{array}[]{ll}\operatorname{sign}(g)&\text{if }r=\infty,\\ g/\|g\|_{2}&\text{if }r=2,\\ \operatorname{sign}({g}_{k^{\prime}})\bm{e}_{k^{\prime}}\text{ with }k^{\prime}\in\arg\max_{k}|g_{k}|&\text{if }r=1,\end{array}\right$. (1)
and projection $\Pi_{r,{x},\kappa\epsilon}$ by
$\Pi_{r,{x},R}\colon x\mapsto\arg\min_{\xi}\left\{\|\xi-x\|_{2}^{2}\mid\|\xi\|_{r}\leq R\right\}.$ (2)
Wasserstein Distributionally Robust Optimization (WDRO)
Robustness guarantees and certificates aim to make model predictions trustworthy under adversarial manipulation (Wong & Kolter, 2018; Cohen et al., 2019; Salman et al., 2019). Empirical risk minimization model $\inf_{\theta}\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]$ optimizes average performance on the observed data but offers no protection against worst-case shifts nearby. Distributionally robust optimization (DRO) addresses this by choosing parameters that perform well against all distributions within a prescribed neighborhood: $\inf_{\theta}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)].$ Here, the worst-case loss is taken over all admissible distributions $\mathbb{P}\in\mathcal{P}$. The ambiguity (or uncertainty) set $\mathcal{P}$ is often constructed by collecting all distributions $\mathbb{P}$ that are similar to the empirical distribution $\mathbb{P}_{N}$.
In this work, we focus on the Wasserstein ambiguity set, which is a ball centered at $\mathbb{P}_{N}$ under the Wasserstein distance. Given a ground cost $d$ on the space of data $\mathcal{Z}$, the Wasserstein distance (Villani, 2008) between two distributions $\mathbb{P}$ and $\mathbb{Q}$ it is defined as $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\left(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{Z}\times\mathcal{Z}}d^{p}(z^{\prime},z)\,\mathrm{d}\pi(z^{\prime},z)\right)^{1/p}\text{ for}\,p\in[1,\infty);$, and $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\operatorname{ess.sup}_{{\pi}}(d)\text{ for}\,p=\infty$. Intuitively, the Wasserstein distance between two distributions $\mathbb{P}$ and $\mathbb{Q}$ is defined as the minimum cost to transport the mass of $\mathbb{P}$ to $\mathbb{Q}$. The WDRO problem with a given budget of perturbation $\epsilon>0$ can be written as
$\inf_{\theta}\sup_{\mathbb{P}\in\Omega_{p}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\text{ where }\Omega_{p}=\left\{\mathbb{P}\mid\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\right\}.$ (3)
It is worth noting that $\mathcal{W}_{d,p}\leq\mathcal{W}_{d,p^{\prime}}$ if $p\geq p^{\prime}$, thus $\Omega_{1}\supseteq\Omega_{p}\supseteq\Omega_{p^{\prime}}\supseteq\Omega_{\infty}$ (see Figure 1).
Lipschitz Certificate
For $p=1$, the worst-case risk over a Wasserstein ball admits the standard Lipschitz upper bound
$\sup_{\mathbb{P}\in\Omega_{1}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\leq\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+L\epsilon.$ (4)
where $L$ is any Lipschitz constant of $z\mapsto\ell(z;\theta)$ with respect to the ground cost. This inequality follows from weak duality and is widely used to make the WDRO objective tractable: one replaces the inner maximization by the surrogate $L\epsilon$ and then controls $L$ (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024). In practice, estimating $L$ reduces to bounding the network’s (global or local) Lipschitz modulus, e.g., fast global products of per-layer operator norms (Virmaux & Scaman, 2018) or tighter activation-aware/local certificates (Jordan & Dimakis, 2020; Shi et al., 2022).
Adversarial Attack.
Adversarial attack methods often construct a perturbed distribution by shifting each sample $X^{(i)}$ along a specific adversarial direction $u^{(i)}$ to get $X_{\mathrm{adv}}^{(i)}$ (Goodfellow et al., 2014; Moosavi-Dezfooli et al., 2016; Carlini & Wagner, 2017). These methods are essentially point-wise attacks, which draws a distribution $\mathbb{P}_{\mathrm{adv}}=\sum_{i=1}^{N}\frac{1}{N}\bm{\delta}_{X^{(i)}_{\mathrm{adv}}}$ in the Wasserstein ambiguity set $\Omega_{p}=\{\mathbb{P}\colon\mathcal{W}_{d,p}\leq\epsilon\}$ when $p=\infty$ (see Figure 1). Whereas, in the $p=1$ case, the ambiguity set only constrains the average transportation cost under an optimal coupling. Hence, the adversary may move some points farther and others less as long as the mean cost stays within budget. This creates a significant gap between the robustness measured against $\Omega_{p=\infty}$ attacks and the theoretical robustness or Lipschitz certificates 4 which are developed for $\Omega_{p=1}$ (Mohajerin Esfahani & Kuhn, 2018; Carlini et al., 2019; Rice et al., 2021).
3 Tractable Interpretation of WDRO for Neural Networks
For certain shallow and convex models (e.g., linear regression, support vector machines, etc.), the tractable representation of the WDRO problem 3 is well-established in the literature (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024). This tractable form enables a computational advantage and provides a clear interpretation of the robustness of regularization mechanisms. In that line of work, the Lipschitz constant often provides a practical and tight upper bound of the corresponding upper bounds. However, when the loss is non-convex, the Lipschitz certificate is not always tight, as outlined in the following remark.
Remark.
Consider a single-point empirical $\mathbb{P}_{N}=\bm{\delta}_{\{X^{(1)}=2\}}$ and a loss given by
$\ell(x)=\begin{cases}|x|\quad\text{if }|x|\leq 1,\\ \frac{1}{2}|x|+\frac{1}{2}\quad\text{otherwise}.\end{cases}$
Then $\ell$ is Lipschitz with modulus 1, however $\sup_{\Omega_{1}}\mathbb{E}_{\mathbb{P}}[\ell(X)]=\mathbb{E}_{\mathbb{P}_{N}}[\ell(X)]+\frac{\epsilon}{2}\quad\text{for any}\,\epsilon>0.$
As presented in the following sections, our main theoretical results (Theorem 3.1 and 3.3) show that Lipschitz modulus provides a tight upper bound for the WDRO problem 3 for a class of ReLU neural networks and smooth activated neural networks.
3.1 Exact and tractable interpretation of WDRO for ReLU neural networks
For a broad class of ReLU networks, the tight (local) Lipschitz constant can be found exactly via activation patterns. For example, for any $H$-layer ReLU network $\theta(x)=W_{H+1}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{H})$, let
$L_{\theta}=\sup_{x\in\mathcal{X}}\sup_{J\in\partial\theta(x)}\|J\|_{r\rightarrow\tilde{r}}$, (5)
where $J\in\partial\theta(x)$ is a general Jacobian of $\theta$ at $x$, then Jordan & Dimakis (2020, Theorem 1) has shown that $\|\theta(x^{\prime})-\theta(x)\|_{\tilde{r}}\leq L_{\theta}\times\|x^{\prime}-x\|_{r}$ for any $x^{\prime},x\in\mathcal{X}$. Moreover, if $\theta$ is in general position (Jordan & Dimakis, 2020, Definition 4), then the chain rule applies and any general Jacobian $J$ must has a form as $W_{H+1}D_{H}W_{H}\cdots D_{1}W_{1}$ for some $[0,1]$-diagonal matrix $D_{h}$, $h=1,\dots,H$. It is worth noting that the set of ReLU networks not in general position is negligible (Jordan & Dimakis, 2020, Theorem 3). Now in equation 5, the maximizer of a convex function (norm operator) must happen at vertices, thus we only need to consider $0/1$-diagonal matrix $D_{h}$.
We formally introduce the concept of mask as follows.
Definition 3.1 (Mask and Cell).
Let $\theta(x)=W_{H+1}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{H})$ be a ReLU network which is in general position. For any tuple $\bm{D}=(D_{1},\dots,D_{H})$, we define
$J_{\bm{D}}=W_{H+1}D_{H}W_{H}\cdots D_{1}W_{1}.$
For any $x\in\mathcal{X}$, we define the set of all 0/1-diagonal masks at $x$ as
$\mathcal{D}_{x}=\left\{\bm{D}=(D_{1},\dots,D_{H})\mid J_{\bm{D}}\in\partial\theta(x),\,D_{h}\text{ is 0/1-diagonal},\,h=1,\dots,H\right\}$
We denote $\mathcal{D}_{\mathcal{X}}=\cup_{x\in\mathcal{X}}\mathcal{D}_{x}$ as the (finite) set of all possible masks.
For any mask $\bm{D}=(D_{1},\dots,D_{H})\in\mathcal{D}_{x}$, let $\mathcal{C}_{\bm{D}}$ be the cell, which is an open linear region, defined by
$\mathcal{C}_{\bm{D}}=\left\{x\mid\operatorname{pre}_{h}(x)_{j}>0\text{ if }D_{h}(j,j)=1\text{ and }\operatorname{pre}_{h}(x)_{j}<0\text{ if }D_{h}(j,j)=0,\,h=1,\dots,H\right\},$
where the pre-activation functions are defined as
$\operatorname{pre}_{h}\colon x\mapsto W_{h}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{h}).$
Given this definition of mask and note that $\mathcal{D}_{\mathcal{X}}$ is finite, one can rewrite equation 5 as $L_{\theta}=\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\|J_{\bm{D}}\|_{r\rightarrow\tilde{r}}$. We adopt this notion and show that it induces an upper bound for the Wasserstein distributional robust optimization (WDRO) problem 3 with cross-entropy loss. Moreover, this upper bound is tight for a class of monotonic ReLU networks.
Theorem 3.1 (WDRO for ReLU).
Given a ReLU network $\theta(x)=W_{H+1}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{H})$ being in general position, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
$\bm{L}\triangleq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}},$ (6)
and
$\bm{l}\triangleq\max_{\begin{subarray}{c}x\in\mathcal{X},\\ \bm{D}\in\mathcal{D}_{x}\end{subarray}}\max_{k^{\prime}\neq k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k^{\prime}}-\bm{e}_{k})^{\top}J_{\bm{D}}u\mid u\in\mathrm{rec}({\mathcal{C}_{\bm{D}}})\right\}.$ (7)
where $J_{\bm{D}}$, $\mathcal{C}_{\bm{D}}$, $\mathcal{D}_{\mathcal{X}}$ are defined in Definition 3.1 and $\mathrm{rec}({\mathcal{C}_{\bm{D}}})$ is the recession cone of $\mathcal{C}_{\bm{D}}$. Then for any $\epsilon>0$, we have
$\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon\leq\sup_{\mathbb{P}\colon\mathcal{W}_{d,1}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\leq\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{L}\epsilon.$ (8)
Moreover, if the dual-norm maximizer $\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of 7, and $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})$ is the largest increment direction of $J_{\bm{D}^{\star}}$, then $\bm{l}=\bm{L}$.
Proof.
To prove inequality 8, we show that $\ell(\cdot,\theta)$ is $\bm{L}$-Lipschitz, and a direction $u$ found in equation 7 induces an admissible attack $\mathbb{P}_{\mathrm{adv}}$ satisfying that $\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$ and $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$. To verify the sufficient condition of $\bm{l}=\bm{L}$, we show that the constructed $\mathbb{P}_{\mathrm{adv}}$ provides $\bm{l}=\bm{L}$. We provide detailed proof in Appendix A.1. ∎
In Figure 2, we illustrate an instance in which our lower and upper bounds match. While equation 7 provides a tight lower bound of the WDRO, it is impractical to scan through all $x\in\mathcal{X}$ and its mask $\mathcal{D}_{x}$. We then introduce a practical lower bound of which we consider the mask associated with the sample points only.
Figure 2: WDRO bounds and PGD attack loss for a fixed $n=K=2$ ReLU classifier with one hidden layer of dimension 8. Lower-bound curves are the cumulative $\bm{l}$ as more reachable activation masks are considered.
Corollary 3.2 (Practical lower bound).
Given assumptions and notations used in Theorem 3.1, let $\mathcal{Z}_{N}=\{(X^{(1)},Y^{(1)}),\dots,(X^{(N)},Y^{(N)})\}$ and
$\bm{l}_{N}\triangleq\max_{\begin{subarray}{c}(X^{(i)},Y^{(i)})\in\mathcal{Z}_{N},\\ \bm{D}\in\mathcal{D}_{x}\end{subarray}}\max_{k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k}-Y^{(i)})^{\top}J_{\bm{D}}u\mid u\in\mathrm{int}(\mathrm{rec}({\mathcal{C}_{\bm{D}}}))\right\}.$ (9)
Then $\bm{l}_{N}\leq\bm{l}$.
In the spirit of formulation 9, we propose a practical adversarial attack in Section 4 which aims to find adversarial direction $u^{(i)}$ for each sample $i$ so that it maximizes the change of the logit function.
3.2 Exact and tractable interpretation of WDRO for smooth activation neural networks
For networks with smooth activations, e.g, GELU (Hendrycks & Gimpel, 2016), SiLU/Swish (Ramachandran et al., 2017; Elfwing et al., 2018), WDRO duality connects worst-case (adversarial) risk to first-order geometry via the Jacobian of the logit map, yielding global Lipschitz-type upper penalties of the form $\sup_{x\in\mathcal{X}}\|J(x)\|_{r\to s}$. Compared to piecewise-linear ReLU certificates, smooth nets trade exact cell-wise constancy for differentiability along rays and curves, suggesting bounds driven by asymptotic Jacobian behavior rather than activation masks.
Let $t$ be a positive scalar, $\theta:\mathcal{X}\subseteq\mathbb{R}^{n}\to\mathbb{R}^{K}$ be a classifier with smooth activations and cross-entropy loss; let $J(x)\in\mathbb{R}^{K\times n}$ be its Jacobian. We then have the following result.
Theorem 3.3 (WDRO for Smooth Networks).
Let $\theta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$ be a differentiable network, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
$\bm{L}\triangleq{2^{1/{s}}}\sup_{x\in{\mathcal{X}}}\left\|\nabla_{x}\theta(x)\right\|_{{r}\to{s}},$
and
$\bm{l}\triangleq\sup_{x\in{\mathcal{X}}}\max_{k^{\prime}\neq k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k^{\prime}}-\bm{e}_{k})^{\top}\nabla_{x}\theta(x)u\right\}.$
Then for any $\epsilon>0$,
$\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon\leq\sup_{\mathbb{P}\colon\mathcal{W}_{d,1}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\leq\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{L}\epsilon.$
In this setting, first-order WDRO penalties are controlled by how $J(x)$ amplifies unit directions and how that amplification projects onto the most competitive non-true logit. The upper slope $\bm{L}$ is the global worst-case amplification, while the lower, margin–directional slope $\bm{l}$ follows rays $x^{(i)}+tu$ and harvests only the component along $(\bm{e}_{k^{\prime}}-\bm{e}_{k_{i}})$. When a ray both attains the global operator norm and aligns with a margin difference, the bound is tight to first order ($\bm{l}=\bm{L}$). This motivates the adversarial procedure used our adversarial attack algorithm WDA (Algogirthm 1), where we search for a direction $u$ and a rival class $k^{\prime}$ that maximize the first-order increase $(\bm{e}_{k^{\prime}}-\bm{e}_{k_{i}})^{\top}J(x)u$.
4 Wasserstein Distributional Attack
Existing point-wise attacks such as FGSM (Goodfellow et al., 2014), DeepFool (Moosavi-Dezfooli et al., 2016), AA (Croce & Hein, 2020), AAA (Liu et al., 2022), keep the adversarial distribution supported on exactly $N$ points, where each point $X^{(i)}_{\operatorname{adv}}$ is perturbed to be precisely on the boundary of the $\epsilon$-ball centered at $X^{(i)}$. To address this issue, we propose a novel method called Wasserstein Distributional Attack (WDA). At a high level, WDA constructs an adversarial distribution, $\mathbb{P}_{\mathrm{adv}}$, supported on a set of $2N$ points. This set comprises of $N$ original empirical samples $X^{(i)}$ and $N$ corresponding adversarial points $X^{(i)}_{\mathrm{adv}}$, each perturbed to an $r$-norm distance $\kappa\epsilon$ from $X^{(i)}$ using the first-order, margin-aligned directions predicted by Theorems 3.13.3, for some $\kappa\geq 1$. In other words,
$\mathbb{P}_{\mathrm{adv}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa}\bm{\delta}_{(X_{\mathrm{adv}}^{(i)},Y^{(i)})}.$ (10)
In the special case where $\kappa=1$, our proposed attack reduces to existing point-wise methods. When $\kappa=2$, WDA simplifies to a uniform distribution over all $2N$ points, with each point receiving a weight of $\frac{1}{2N}$. This $2N$-support mixture belongs to the $\Omega_{1}$ ambiguity set and serves as a constructive, distributional adversary; it is not necessarily the inner maximizer of WDRO. We now make the first-order ascent directions explicit; this is the step used by WDA to realize the margin-aligned rays from Theorems 3.13.3.
Define (sub)gradient $g_{j}(x)\in\partial_{x}\big(\theta_{j}-\theta_{k}\big)(x)$. Then $\mathcal{M}_{r}(g_{j})$ give the per–iteration, first–order version of the ray ascent used in Theorems 3.13.3: within a ReLU cell (affine logits) or for smooth activations (continuous $J$), moving along $u_{j}=\mathcal{M}_{r}(g_{j})$ increases the gap at rate $\|g_{j}(x)\|_{s}$. During an initial probing phase, we evaluate all rivals $j\neq k$ using these first-order steps. At the end of that phase, we fix a single rival $j^{*}$ based on the logits magnitude and continue the remaining iterations. If we allow $j^{*}$ to change at every step, the update can oscillate across classes and chase locally steep but globally suboptimal directions for misclassifications. Finally, we project each step to the ball of radius $\kappa\varepsilon$ around the anchor $X^{(i)}$ to the WDRO budget. The procedure for implementing the Wasserstein Distributional Attack is presented in Algorithm 1. A visualization of our algorithm is shown in Figure 3.
Algorithm 1 Wasserstein Distributional Attack (WDA)
1:Inputs: neural network $\theta:\mathbb{R}^{n}\to\mathbb{R}^{K}$, empirical distribution $\mathbb{P}_{N}=\sum_{i=1}^{N}\bm{\delta}_{(X^{(i)},Y^{(i)})}$, budget $\epsilon>0$, cost-norm $r\in\{1,2,\infty\}$, WDA paramter $\kappa\geq 1$, step size $\alpha>0$, and $0<$ prob $\leq$ maxiter
2:Outputs: Wasserstein distributional attack $\mathbb{P}_{\mathrm{adv}}$ such that $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$ where $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$
3:Initialize: dual-norm maximizer $\mathcal{M}$ (1), projection $\Pi$ (2)
4:for $i=1$ to NN do
5:  $x_{0}\leftarrow X^{(i)}$, $\bm{e}_{k}\leftarrow Y^{(i)}$ for some $k=1,\dots,K$
6:  for iter =0=0 to maxiter do
7:   if iter << prob then $\mathcal{J}=\{1,\dots,K\}\setminus\{k\}$ else $\mathcal{J}=\{j^{*}\}$
8:   $g_{j}\leftarrow\nabla_{x}\theta\left(x_{\mathrm{iter}}\right)^{\top}(\bm{e}_{j}-\bm{e}_{k})$ for $j\in\mathcal{J}$
9:   $u_{j}\leftarrow\mathcal{M}_{r}(g_{j})$ for $j\in\mathcal{J}$
10:   $\varphi_{j}\leftarrow\Pi_{r,X^{(i)},\kappa\epsilon}\left(x_{\mathrm{iter}}+\alpha u_{j}\right)$ for $j\in\mathcal{J}$
11:   $j^{*}=\arg\max_{j\in\mathcal{J}}\theta_{j}\left(\varphi_{j}\right)$
12:   $x_{\mathrm{iter+1}}\leftarrow\varphi_{j^{*}}$
13:  end for
14:  $X^{(i)}_{\mathrm{adv}}\leftarrow x_{\mathrm{maxiter}}$
15:end for
[…]
17:return $\mathbb{P}_{\mathrm{adv}}$
Figure 3: Wasserstein Distributional Attack (WDA, Alg. 1) for $r=2$. At each iteration $x_{t}$, WDA forms K1K\!-\!1 candidates $\varphi_{j}$ and updates using the one with the largest logit $\theta_{j}(\varphi_{j})$. For reference, PGD follows the dual-norm gradient direction; DeepFool linearizes the decision boundary.
5 Related Work
Robustness Certificates
Early scalable global certificates control the Lipschitz constant by multiplying per-layer operator norms, which is fast to compute yet data-agnostic and typically loose on deep nets (Virmaux & Scaman, 2018). For ReLU networks, local (activation-aware) methods exploit piecewise linearity to produce much tighter, input-conditioned certificates on individual activation regions (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018; Shi et al., 2022). Most relevant to exact local Lipschitzness, Jordan & Dimakis (2020) showed that for a broad class of ReLU networks in general position, the local Lipschitz constant can be computed exactly by optimizing over activation patterns.
Adversarial Attacks
Adversarial Attack methods seek for perturbation $x^{\prime}$ formed by adding a small, human-imperceptible perturbation to a clean input $x$ that causes misclassification (Szegedy et al., 2014). The threat model specifies the attacker’s knowledge (white-box vs. black-box), the admissible perturbation set (e.g., $r_{2}$ balls with budget $\epsilon$), and the objective (e.g., worst-case loss within the ball). Canonical white-box methods include FGSM (Goodfellow et al., 2014), multi-step PGD (Madry et al., 2018), CW (Carlini & Wagner, 2017), and gradient-based margin attacks such as DeepFool (Moosavi-Dezfooli et al., 2016). Decision-based and score-free attacks (black-box) include Boundary Attack (Brendel et al., 2021) and Square Attack (Andriushchenko et al., 2020). Robust evaluation is subtle: gradient masking can inflate apparent robustness if attacks are not adapted (Athalye et al., 2018).
To standardize evaluation, AutoAttack (AA) (Croce & Hein, 2020) composes strong, parameter-free attacks (APGD-CE, APGD-DLR, FAB, Square) and is widely adopted for reporting robust accuracy. RobustBench (Croce et al., 2021) curates model zoos and standardized test protocols across datasets and $r_{p}$ threat models, enabling comparable and reproducible robustness claims. Liu et al. (2022) proposed Adaptive Auto Attack (A3), which incorporates Adaptive Direction Initialization (ADI) and Online Statistics-based Discarding (ODS) (Tashiro et al., 2020) to enhance attack efficiency. In our experiments, we report robustness under AA and A3 following RobustBench conventions and use them as baselines for comparison.
Several works have focused on adversarial attacks tailored to ReLU networks. Croce & Hein (2018) introduced rLR-QP, a gradient-free method that navigates the piecewise-linear regions of ReLU models by solving convex subproblems and enhancing exploration with randomization and local search. More recently, Zhang et al. (2022) developed BaB-Attack, a branch-and-bound framework that operates in activation space, leveraging bound propagation, beam search, and large neighborhood search to uncover stronger adversarial examples than conventional gradient-based approaches, particularly on hard-to-attack inputs.
6 Experiments
Experimental settings
To evaluate the effectiveness of WDA, we test the adversarial robustness of several state-of-the-art defense models on the CIFAR-10 dataset. For comparison, we report both $r_{\infty}$ and $r_{2}$ robustness under perturbation budgets of $\epsilon=8/255$ and $\epsilon=0.5$, respectively. In addition to point-wise attack, we conduct a separate Wasserstein distributional attack experiment to further validate our method. Specifically, we set $\kappa=2$ in Algorithm 1 to find the adversarial (distributional) attack $\mathbb{P}_{\mathrm{adv}}$equation 10 and reporting classification accuracy on the distribution by $(1-1/\kappa)\times\mathrm{cleanaccuracy}+\kappa\times\mathrm{adversarialaccuracy}$. Our attack is benchmarked against AA, APGD-DLR, APGD-CE (Croce & Hein, 2020), and A3 (Liu et al., 2022). The defense models, along with their official implementations and pretrained weights, are obtained from RobustBench Croce et al. (2021). All experiments are conducted on 2x NVIDIA GeForce RTX 4090 GPU and 1x NVIDIA H200.
6.1 Comparison with existing baselines
Setup
For the baselines AA, APGD-DLR, APGD-CE, and A3, we adopt the configurations reported in their respective research papers. Meanwhile, in WDA, we set the number of probe steps to $\alpha_{probe}=10$, and use $\alpha_{atk}=20$ attack iterations. We further propose A3++, an extension of A3 that incorporates our attack into its framework.
Table 1: Comparison of robust accuracy of WDA and baseline methods against various defenses on CIFAR-10 with $r_{\infty}$ perturbations ($\epsilon=8/255$) and $r_{2}$ perturbations ($\epsilon=0.5$), as well as on CIFAR-100 with $r_{\infty}$ perturbations ($\epsilon=8/255$). The first column reports the source paper corresponding to each defense method.
Paper Model Clean Single method Ensemble method
APGD -CE APGD -DLR WDA ($\kappa=1$) WDA ($\kappa=2$) AA A3 A3++
[…]
Bartoldson et al. (2024) WRN-94-16 93.68 76.15 74.31 74.05 65.25 73.71 73.55 73.54
Bartoldson et al. (2024) WRN-82-8 93.11 74.17 72.54 71.85 62.06 71.59 71.46 71.46
Cui et al. (2024) WRN-28-10 92.16 70.60 68.62 68.07 60.01 67.73 67.58 67.57
Wang et al. (2023) WRN-70-16 93.25 73.46 71.68 71.02 63.08 70.69 70.53 70.52
Wang et al. (2023) WRN-28-10 92.44 70.24 68.24 67.60 60.96 67.31 67.17 67.17
Xu et al. (2023) WRN-28-10 93.69 67.08 69.00 66.39 63.25 63.89 63.93 63.84
Sehwag et al. (2022) RN-18 84.59 58.40 57.66 56.30 54.65 55.54 55.50 55.50
[…]
Wang et al. (2023) WRN-70-16 95.54 85.66 85.30 85.00 77.63 84.97 84.96 84.97
Wang et al. (2023) WRN-28-10 95.16 84.52 83.89 83.71 76.31 83.68 83.68 83.68
Sehwag et al. (2022) WRN-34-10 90.93 78.23 78.16 77.51 72.01 77.24 77.22 77.25
Sehwag et al. (2022) RN-18 89.76 75.24 75.32 74.69 69.75 74.41 74.41 74.40
Ding et al. (2020) WRN-28-4 88.02 66.62 66.62 66.22 63.04 66.09 66.05 66.06
Cui et al. (2024) WRN-28-10 89.05 66.58 67.08 66.59 64.19 66.44 66.41 66.42
[…]
Wang et al. (2023) WRN-28-10 72.58 44.09 39.66 39.12 43.61 38.77 38.70 38.71
Addepalli et al. (2022) RN-18 65.45 33.47 28.82 28.26 37.64 27.67 27.65 27.63
Cui et al. (2024) WRN-28-10 73.85 43.82 40.37 39.57 43.68 39.18 39.17 39.14
Robustness on $r_{\infty}$ and $l_{2}$ Perturbations for CIFAR-10
Table 1 presents the robust accuracy of several attack methods under $r_{\infty}$ perturbations with $\epsilon=8/255$ and $r_{2}$ perturbations with $\epsilon=0.5$. Across both single and ensemble based threat models, WDA consistently outperforms other single-method attacks (APGD-CE and APGD-DLR), highlighting its effectiveness as a stronger standalone adversarial evaluation. Moreover, WDA achieves results that are often comparable to ensemble-based methods, indicating its ability to match the strength of more computationally demanding attack aggregations. Within the ensemble family, A3++ demonstrates clear improvements over AA and provides competitive performance with A3, surpassing it on several defense models (3 out of 7 under $r_{\infty}$ and 1 out of 6 under $r_{2}$). Notably, under $r_{\infty}$ and $r_{2}$, WDA ($\kappa=2$) produces lower robust accuracy values than other attacking methods across all defense models. This highlights the potential of Wasserstein distributional attack.
6.2 Ablation study
Figure 4: Robust accuracy with varying $\kappa$ on different defense methods.
Figure 5: Convergence of Wang et al. (2023) under $\ell_{2}$ perturbations ($\epsilon=0.5$).
Analysis on the effect of different parameter on robust accuracy
Figure 4 illustrates how the robust accuracy of WDA varies as the parameter $\kappa$ increases. Across all models, raising $\kappa$ beyond 1 generally leads to a noticeable drop in robust accuracy. Specifically, for the models from Wang et al. (2023), Sehwag et al. (2022), and Sehwag et al. (2022) (RN-18), the best performance is observed at $\kappa=3$, whereas for Ding et al. (2020), the lowest robust accuracy occurs at $\kappa=2$. These results indicate that increasing $\kappa$ consistently weakens model robustness, with the precise $\kappa$ that produces the largest drop depending on the architecture. Figure 5 presents robust accuracy for the Wang et al. (2023) model with $r_{2}$ perturbations and with different step sizes for attack. The x-axis represents iterations, and the y-axis shows robust accuracy. In Figure 5, smaller step sizes (0.01 and 0.005) lead to higher robust accuracy (97%–99%), reflecting weaker attacks. The most effective attack occurs at step size 0.2, where the accuracy drops to around 88%. At larger step sizes (0.5 and 1.0), robust accuracy stabilizes at lower values despite initial drops, suggesting reduced attack effectiveness.
7 Conclusions
We presented tight robustness certificates and stronger adversarial attacks for deep neural networks by exploiting their local geometric structure. For ReLU networks, we derived exact WDRO bounds using their piecewise-affine property, computing data-dependent Lipschitz constants from activation patterns that significantly tighten existing global bounds. For networks with smooth activations (GELU, SiLU), we characterized the worst-case loss through asymptotic Jacobian behavior along adversarial rays, providing the first tractable WDRO analysis for these modern architectures. Our Wasserstein Distributional Attack (WDA) algorithm constructs adversarial distributions on $2N$ points rather than restricting to $N$ perturbed points, achieving lower robust accuracy than state-of-the-art methods across CIFAR-10/100 benchmarks. While WDA incurs additional computational overhead compared to single-point attacks due to evaluating multiple candidate perturbations per iteration, it demonstrates that existing robustness evaluations significantly underestimate vulnerability by considering only point-wise perturbations. Together, these contributions narrow the gap between theoretical certificates and practical evaluation, revealing that both tighter bounds and stronger attacks emerge from properly leveraging network geometry and distributional perspectives.
Appendix A Proofs
A.1 Proof of Theorem 3.1
Proof of Upper Bound
It is a standard result that for any $y=\bm{e}_{k}$, if $\ell$ is the cross-entropy loss then
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\log\left[\operatorname{softmax}\theta(x^{\prime})\right]_{k}-\log\left[\operatorname{softmax}\theta(x)\right]_{k}\right|\\ &\leq 2^{1/s}\|\theta(x^{\prime})-\theta(x)\|_{s},\end{array}$ (11)
or if $\ell$ is the DLR loss then
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\left(\max_{k_{1}\neq k}\theta(x^{\prime})_{k_{1}}-\theta(x^{\prime})_{k}\right)-\left(\max_{k_{2}\neq k}\theta(x)_{k_{2}}-\theta(x)_{k}\right)\right|\\ &\leq 2^{1/s}\|\theta(x^{\prime})-\theta(x)\|_{s}.\end{array}$ (12)
In addition, by Jordan & Dimakis (2020), we have that for any $x^{\prime},x\in\mathcal{X}$,
$\|\theta(x^{\prime})-\theta(x)\|_{s}\leq\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}\times\|x^{\prime}-x\|_{r}.$ (13)
Thus,
$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&\leq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}\times\|x^{\prime}-x\|_{r}\\ &=\bm{L}\times d((x^{\prime},y),(x,y)).\end{array}$ (14)
for any $x^{\prime},x\in\mathcal{X}$ and therefore by using Lipschitz certificate (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024; Chu et al., 2024), we have
$\sup_{\mathbb{P}\colon\mathcal{W}_{d,1}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\leq\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{L}\epsilon,$ (15)
for any $\epsilon>0$.
Proof of Lower Bound
To show that the lower bound of the worst-case loss is $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$, it is enough to construct a perturbation $\tilde{Z}$, a weight $\eta\in(0,1]$, and a distribution
$\mathbb{P}_{\mathrm{adv}}=\sum_{i=1,i\neq\iota}^{N}\frac{1}{N}\bm{\delta}_{Z^{(i)}}+\frac{1-\eta}{N}\bm{\delta}_{Z^{(\iota)}}+\frac{\eta}{N}\bm{\delta}_{\tilde{Z}}$, (16)
so that $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$ and $\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$.
Since $\mathcal{D}_{\mathcal{X}}$ is finite, let $x^{\star},\bm{D}^{\star},k^{\prime\star},k^{\star}$ and sequence $\{u^{\star}_{t}\}$ be the maximizer in 7, i.e., $\bm{D}^{\star}\in\mathcal{D}_{x^{\star}},k^{\prime\star}\neq k^{\star},\{u^{\star}_{t}\}\subset\operatorname{int}(\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}}))$ and
$(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})^{\top}J_{\bm{D}^{\star}}u^{\star}_{t}\rightarrow\bm{l}\text{ when }t\rightarrow\infty.$
In particular, $\theta$ is affine and differentiable on $\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}})$. Since $u^{\star}_{t}$ belongs to the open cone $\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}})$, one has that for any $\alpha>0$,
$\tilde{x}=x^{\star}+\alpha u^{\star}_{t}\in\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}}),$ (17)
and thus
$\nabla_{x}\theta(\tilde{x})=J_{\bm{D^{\star}}},\quad\theta(\tilde{x})-\theta(x^{\star})=\alpha J_{\bm{D^{\star}}}u^{\star}_{t}.$ (18)
Choose root $\iota$ so that $(X^{(\iota)},Y^{(\iota)}=\bm{e}_{k^{\star}})$. Then when $\ell$ is the cross-entropy loss or DLR loss, by a technical Lemma A.1 one has
$\lim_{\alpha\to\infty}\frac{\Delta\ell(\alpha)}{\alpha}=\lim_{\alpha\to\infty}\frac{\ell(\tilde{x},Y^{(\iota)};\theta)-\ell(x^{\star},Y^{(\iota)};\theta)}{\alpha}\geq v_{k^{\prime\star}}-v_{k^{\star}}.$ (19)
where $v=J_{\bm{D}^{\star}}u^{\star}_{t}$. Now choose $\alpha$ large enough so that $\Delta\ell(\alpha)\approx\alpha(v_{k^{\prime\star}}-v_{k^{\star}})$, $\Delta\ell(\alpha)\gg\ell(x^{\star},Y^{(\iota)};\theta)-\ell(X^{(\iota)},Y^{(\iota)};\theta)$, and $N\epsilon<\|\tilde{x}-X^{(i)}\|_{r}\approx\alpha$. Set $\tilde{Z}=(\tilde{x},Y^{(\iota)})$, then
$\begin{array}[]{rl}\ell(\tilde{Z};\theta)-\ell(Z^{(\iota)};\theta)&=\Delta\ell(\alpha)+\ell(x^{\star},Y^{(\iota)};\theta)-\ell(X^{(\iota)},Y^{(\iota)};\theta)\\ &\approx\|\tilde{x}-X^{(i)}\|_{r}(v_{k^{\prime\star}}-v_{k^{\star}})\\ &=d(\tilde{Z},Z^{(\iota)})\times(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})^{\top}J_{\bm{D}^{\star}}u^{\star}_{t}\xrightarrow{t\rightarrow\infty}\bm{l}\times d(\tilde{Z},Z^{(\iota)}).\end{array}$ (20)
Now set $\eta=\frac{N\epsilon}{d(\tilde{Z},Z^{(\iota)})}\in(0,1]$, then
$\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\frac{\eta}{N}d(\tilde{Z},Z^{(\iota)})=\epsilon.$ (21)
Moreover,
$\begin{array}[]{rl}\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]&=\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\frac{\eta}{N}\left(\ell(\tilde{Z};\theta)-\ell(Z^{(\iota)};\theta)\right)\\ &\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\frac{\epsilon}{d(\tilde{Z},Z^{(\iota)})}\bm{l}d(\tilde{Z},Z^{(\iota)})\\ &=\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon.\end{array}$ (22)
Therefore, the lower bound of the worst-case loss is $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$.
Sufficient condition of $\bm{l}=\bm{L}$
Suppose that the dual-norm maximizer $\xi=\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of 7, then we have
$\bm{l}$ $=(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})^{\top}J_{\bm{D}^{\star}}u^{\star}_{t}$ (23)
$\geq(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})^{\top}J_{\bm{D}^{\star}}\xi$ $\text{ (since }u^{\star}_{t}\text{ is the maximizer)}$
$=\|(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})^{\top}J_{\bm{D}^{\star}}\|_{s}$ (by definition of dual-norm maximizer)
$=\|(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})\|_{r}\|J_{\bm{D}^{\star}}\|_{r\rightarrow s}$
$=2^{1/s}\|J_{\bm{D}^{\star}}\|_{r\rightarrow s}=\bm{L},$
where the second last equality holds true because $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})$ is the largest increment direction of $J_{\bm{D}^{\star}}$.
Figure 6: Illustration of Proof of Lower Bound
A.2 Proof of Theorem 3.3
Proof.
The proof for the upper and lower bounds is similar to the methodology we discussed in our previous exchange.
Proof of Upper Bound The WDRO upper bound is a direct consequence of the Lipschitz continuity of the loss function. The Lipschitz constant of the combined loss function, $L_{\ell}=\sup_{Z\in\mathcal{Z}}\|\nabla_{x}\ell(Z;\theta)\|_{r}$, is bounded by the product of the Lipschitz constant of the loss with respect to the output and the Lipschitz constant of the network. That is,
$\bm{L}\leq\|J(x)\|_{r\to s}\cdot\|\nabla_{\theta}\ell\|_{s}=\|J(x)\|_{r\to s}\cdot 2^{1/s}$
Proof of Lower Bound The proof of the lower bound is identical with the ReLU network case, where it relies on constructing a specific adversarial distribution. This finds a point $x^{\star}$ and a direction $u^{\star}$ that maximize the rate of change of the loss. The constant $\bm{l}$ is defined as this maximum rate of change. By constructing a perturbed point $\tilde{x}=x^{\star}+\alpha u^{\star}$ and a corresponding adversarial distribution, it is shown that the worst-case loss is at least $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$. ∎
A.3 Technical Proofs
Lemma A.1 (Technical lemma).
In equation 19, if $\ell=\ell_{CE}$ is the cross-entropy loss, then
$\lim_{\alpha\to\infty}\frac{\Delta\ell_{CE}}{\alpha}=\max_{i}(J_{\bm{D}^{\star}}u^{\star}_{t})_{i}-(J_{\bm{D}^{\star}}u^{\star}_{t})_{k^{\star}}.$
Else if $\ell=\ell_{DLR}$ is the DLR loss, then
$\lim_{\alpha\to\infty}\frac{\Delta\ell_{DLR}}{\alpha}=\max_{i\neq k^{\star}}(J_{\bm{D}^{\star}}u^{\star}_{t})_{i}-(J_{\bm{D}^{\star}}u^{\star}_{t})_{k^{\star}}.$
Proof.
Let $\theta^{\star}=\theta(x^{\star})$ and the change in network output be $\Delta\theta=\theta(\tilde{x})-\theta(x^{\star})=\alpha J_{\bm{D}^{\star}}u^{\star}_{t}$. We will analyze the limit for each loss function separately.
Cross-Entropy Loss: The difference in loss is $\Delta\ell_{CE}=\ell_{CE}(\theta(\tilde{x}),e_{k^{\star}})-\ell_{CE}(\theta(x^{\star}),e_{k^{\star}})$. Using the property $\ell_{CE}(z,e_{k^{\star}})=-(z_{k^{\star}}-\log\sum_{k}e^{z_{k}})$, the loss difference is:
$\Delta\ell_{CE}=-\Delta\theta_{k^{\star}}+\log\left(\sum_{k}e^{\Delta\theta_{k}}\cdot\operatorname{softmax}(\theta^{\star})_{k}\right)$
To find the limit of the average rate of change, $\frac{\Delta\ell_{CE}}{\alpha}$, we substitute $\Delta\theta=\alpha v$, where $v_{k}=(J_{\bm{D}^{\star}}u^{\star}_{t})_{k}$.
$\lim_{\alpha\to\infty}\frac{\Delta\ell_{CE}}{\alpha}=\lim_{\alpha\to\infty}\left[\frac{1}{\alpha}\log\left(\sum_{k}\operatorname{softmax}(\theta^{\star})_{k}e^{\alpha v_{k}}\right)-v_{k^{\star}}\right]$
Let $v_{\max}=\max_{k}v_{k}$. Factoring out the dominant term $e^{\alpha v_{\max}}$ from the sum, the expression becomes:
$\lim_{\alpha\to\infty}\left[\frac{1}{\alpha}\left(\log(e^{\alpha v_{\max}})+\log\left(\sum_{k}\operatorname{softmax}(\theta^{\star})_{k}e^{\alpha(v_{k}-v_{\max})}\right)\right)-v_{k^{\star}}\right]$
$\lim_{\alpha\to\infty}\left[v_{\max}+\frac{1}{\alpha}\log\left(\sum_{k}\operatorname{softmax}(\theta^{\star})_{k}e^{\alpha(v_{k}-v_{\max})}\right)-v_{k^{\star}}\right]$
The sum inside the logarithm converges to a constant value, as all terms with $v_{k}<v_{\max}$ go to 0. The logarithmic term is therefore bounded. The term $\frac{1}{\alpha}$ causes the entire second term to go to 0. The limit is thus:
$v_{\max}-v_{k^{\star}}=\max_{k}(J_{\bm{D}^{\star}}u^{\star}_{t})_{k}-(J_{\bm{D}^{\star}}u^{\star}_{t})_{k^{\star}}$
DLR Loss: The difference in DLR loss is $\Delta\ell_{DLR}=\ell_{DLR}(\theta(\tilde{x}),k^{\star})-\ell_{DLR}(\theta(x^{\star}),k^{\star})$.
$\Delta\ell_{DLR}=\left(\max_{k\neq k^{\star}}\theta(\tilde{x})_{k}-\theta(\tilde{x})_{k^{\star}}\right)-\left(\max_{k\neq k^{\star}}\theta(x^{\star})_{k}-\theta(x^{\star})_{k^{\star}}\right)$
Substituting $\theta(\tilde{x})=\theta^{\star}+\Delta\theta$:
$\Delta\ell_{DLR}=\left(\max_{k\neq k^{\star}}(\theta^{\star}_{k}+\Delta\theta_{k})-\max_{k\neq k^{\star}}\theta^{\star}_{k}\right)-\Delta\theta_{k^{\star}}$
To find the limit of the average rate of change, $\frac{\Delta\ell_{DLR}}{\alpha}$, we substitute $\Delta\theta=\alpha v$ and analyze as $\alpha\to\infty$.
$\lim_{\alpha\to\infty}\frac{\Delta\ell_{DLR}}{\alpha}=\lim_{\alpha\to\infty}\frac{1}{\alpha}\left(\max_{k\neq k^{\star}}(\theta^{\star}_{k}+\alpha v_{k})-\max_{k\neq k^{\star}}\theta^{\star}_{k}\right)-v_{k^{\star}}$
As $\alpha\to\infty$, the term $\alpha v_{k}$ dominates inside the maximum function. The limit of the maximum term is therefore $\max_{k\neq k^{\star}}v_{k}$.
$\lim_{\alpha\to\infty}\frac{\Delta\ell_{DLR}}{\alpha}=\left(\max_{k\neq k^{\star}}v_{k}\right)-v_{k^{\star}}=\max_{k\neq k^{\star}}(J_{\bm{D}^{\star}}u^{\star}_{t})_{k}-(J_{\bm{D}^{\star}}u^{\star}_{t})_{k^{\star}}$
Appendix B Additional result
Figure 7: Comparison of step size effect on model Cui et al. (2024) under $r_{\infty}$ perturbation ($\epsilon=8/255$).
Figure 7 illustrates the convergence behavior of Cui et al. (2024) in terms of robust accuracy. Larger step sizes (0.1, 0.5) lead to higher final accuracy, whereas a smaller step size of 0.02 results in the lowest robust accuracy, indicating the most effective attack.
Figure 8: Comparison of robust accuracy with varying $\kappa$
Figure 8 presents the robust accuracy across different values of $\kappa$. Among the tested settings, $\kappa=2$ consistently produces the lowest robust accuracy for all models (Cui et al. (2024), Wang et al. (2023), and Sehwag et al. (2022)).
Appendix C Usage of Large Language Models (LLMs)
We used ChatGPT solely for revising the writing of the paper. Note that revision here strictly means enhancing the clarity and readability of the text (e.g., fixing typos or constructing latex tables), and not for any other purposes.

### v2

@@
 Wasserstein distributionally robust optimization (WDRO) provides a framework for adversarial robustness, yet existing methods based on global Lipschitz continuity or strong duality often yield loose upper bounds or require prohibitive computation.
-In this work, we address these limitations by introducing a primal approach and adopting a notion of exact Lipschitz certificate to tighten this upper bound of WDRO.
-In addition, we propose a novel Wasserstein distributional attack (WDA) that directly constructs a candidate for the worst-case distribution.
-Compared to existing point-wise attack and its variants, our WDA offers greater flexibility in the number and location of attack points.
-In particular, by leveraging the piecewise-affine structure of ReLU networks on their activation cells, our approach results in an exact tractable characterization of the corresponding WDRO problem.
-Extensive evaluations demonstrate that our method achieves competitive robust accuracy against state-of-the-art baselines while offering tighter certificates than existing methods.
+We address these limitations with a primal approach and adopt a notion of exact Lipschitz certificates to tighten this upper bound of WDRO.
+For ReLU networks, we leverage the piecewise-affine structure on activation cells to obtain an exact tractable characterization of the corresponding WDRO problem.
+We further extend our analysis to modern architectures with smooth activations (e.g., GELU, SiLU), such as Transformers.
+Additionally, we propose novel Wasserstein Distributional Attacks (WDA, WDA++) that construct candidates for the worst-case distribution.
+Compared to existing attacks that are restricted to point-wise perturbations, our methods offer greater flexibility in the number and location of attack points.
+Extensive evaluations demonstrate that our proposed framework achieves competitive robust accuracy against state-of-the-art baselines while offering tighter certificates than existing methods.
 Our code is available at https://github.com/OLab-Repo/WDA
 1 Introduction
-Modern deep networks achieve remarkable accuracy yet remain fragile to distribution shift and adversarial perturbations (Szegedy et al., 2014; Goodfellow et al., 2014; Kurakin et al., 2018; Hendrycks & Dietterich, 2019; Ovadia et al., 2019; Taori et al., 2020; Koh et al., 2021), raising concerns about their reliability in deployment.
-A principled avenue for robustness is Wasserstein distributionally robust optimization (WDRO, Mohajerin Esfahani & Kuhn 2018; Gao & Kleywegt 2023), which controls worst-case test risk over an ambiguity set within a Wasserstein ball around the empirical distribution and admits tight dual characterizations from optimal transport (Villani, 2008; Santambrogio, 2015).
-While numerous defenses have been proposed, a fundamental gap persists between theoretical robustness certificates and practical adversarial evaluation: existing Lipschitz-based certificates often provide loose upper bounds that vastly overestimate the true worst-case loss (Virmaux & Scaman, 2018), while standard attacks restrict perturbations to fixed-radius balls around individual points (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018).
-This mismatch stems from two limitations: certificates typically rely on global worst-case analysis that ignores the actual network geometry traversed by data, and attacks consider only point-wise perturbations rather than distributional shifts permitted by Wasserstein threat models (Singh et al., 2018; Gao & Kleywegt, 2023).
-The discrepancy is particularly pronounced for modern architectures with ReLU activations, where the network behaves as a piecewise-affine function whose local properties vary dramatically across regions (Jordan & Dimakis, 2020), and those with smooth activations (GELU, SiLU/Swish) exhibit complex nonlinear geometry (Hendrycks & Gimpel, 2016; Ramachandran et al., 2017; Elfwing et al., 2018).
+Modern deep networks achieve remarkable accuracy yet remain fragile to distribution shift and adversarial perturbations Szegedy et al.
+(2014); Goodfellow et al.
+(2014); Kurakin et al.
+(2018); Hendrycks and Dietterich (2019); Ovadia et al.
+(2019); Taori et al.
+(2020); Koh et al.
+(2021), raising concerns about their reliability in deployment.
+A principled avenue for robustness is Wasserstein distributionally robust optimization (WDRO, Mohajerin Esfahani and Kuhn 2018; Gao and Kleywegt 2023), which controls worst-case test risk over an ambiguity set within a Wasserstein ball around the empirical distribution and admits tight dual characterizations from optimal transport (Villani and others, 2008; Santambrogio, 2015).
+While numerous defences have been proposed, a fundamental gap persists between theoretical robustness certificates and practical adversarial evaluation: existing Lipschitz-based certificates often provide loose upper bounds that vastly overestimate the true worst-case loss (Virmaux and Scaman, 2018), while standard attacks restrict perturbations to fixed-radius balls around individual points (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018).
+This mismatch stems from two limitations: certificates typically rely on global worst-case analysis that ignores the actual network geometry traversed by data, and attacks consider only point-wise perturbations rather than distributional shifts permitted by Wasserstein threat models (Singh et al., 2018; Gao and Kleywegt, 2023).
+The discrepancy is particularly pronounced for modern architectures with ReLU activations, where the network behaves as a piecewise-affine function whose local properties vary dramatically across regions (Jordan and Dimakis, 2020), and those with smooth activations (GELU, SiLU/Swish) exhibit complex nonlinear geometry (Hendrycks and Gimpel, 2016; Ramachandran et al., 2017; Elfwing et al., 2018).
 In this work, we aim to address both sides of this gap: our contributions can be summarized as follows.
   1. 1.
-    For a class of networks with Rectified Linear Unit (ReLU) activations (Nair & Hinton, 2010), we analyze the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem by connecting with the tight Lipschitz constant studied in Jordan & Dimakis, 2020.
-Our analysis is based on the classical underlying piecewise-affine structure of ReLU networks: on any strict ReLU cell, the logit map $\theta(\cdot)$ is affine with a constant input-logit Jacobian $J_{D}$.
-Our contribution is to leverage this structure for WDRO, which requires combining the Lipschitz constant of the logit map and the sensitivity of the softmax cross-entropy, or the DLR loss.
-Our first theoretical result yields an upper bound of WDRO induced by $\bm{L}\triangleq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}$, where $J_{\bm{D}}$ is general Jacobian of the logit map.
-(See 3.1 for precise definition of $J_{\bm{D}}$.) In addition, we derive a lower bound of WDRO by constructing a concrete and finite worst-case distribution.
-(See equation 16 for the explicit formulation.) This worst-case distribution is constructed by perturbing the empirical sample along the direction which the logit map is most varied.
+    For a class of networks with Rectified Linear Unit (ReLU) activations (Nair and Hinton, 2010), we analyse the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem by connecting with the tight Lipschitz constant studied in Jordan and Dimakis, 2020.
+Our first theoretical result yields an upper bound of WDRO induced by $\bm{L}\triangleq 2^{1/r}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}$, where $J_{\bm{D}}$ is general Jacobian of the logit map, see 3.1.
+In addition, we derive a lower bound of WDRO by constructing a concrete and finite worst-case distribution, see equation 17.
+This worst-case distribution is constructed by perturbing the empirical sample along the direction in which the logit map is most varied.
 Moreover, we provide a sufficient condition where our lower and upper bounds match, and simulate an instance to illustrate this tightness, see Figure 1.
+It is worth noting that not every Lipschitz constant yields a tight bound of WDRO, see Remark Remark.
   2. 2.
-    We further analyze the upper and lower bounds of the Wasserstein Distributional Robust Optimization (WDRO) problem for a class of MLP with smooth activation and cross-entropy loss.
-Unlike ReLU activation or DLR loss, which might create degeneration edges, the chain rule is readily applied in this case, and the Lipschitz constant of the loss is naturally computed by estimating its gradients.
-Similar to the analysis of the ReLU networks, we obtain the upper bound of the WDRO as $\bm{L}\triangleq 2^{1/s}\max_{x\in\mathcal{X}}\left\|\nabla\theta(x)\right\|_{{r}\to{s}}$ while the worst-case distribution and lower bound are constructed similar to the ReLU networks.
+    We further analyse the upper and lower bounds of the WDRO problem for a class of architectures with smooth activation and cross-entropy loss.
+Unlike ReLU activation or DLR loss, which might create degeneration edges, the global differentiability of smooth networks enables an exact characterization of the Lipschitz constant via the gradient norm.
+Similar to the analysis of the ReLU networks, we obtain the upper bound of the WDRO as $\bm{L}\triangleq 2^{1/r}\max_{x\in\mathcal{X}}\left\|\nabla\theta(x)\right\|_{{r}\to{s}}$ while the worst-case distribution and lower bound are constructed similar to the ReLU networks.
   3. 3.
-    Finally, we bridge the gap between WDRO theory and adversarial evaluation by introducing the Wasserstein Distributional Attack (WDA), which directly constructs adversarial distributions within the Wasserstein ball rather than restricting to point-wise perturbations.
-Unlike existing attacks that place all adversarial examples on the $\epsilon$-ball boundary, WDA flexibly interpolates between point-wise ($\kappa=1$) and truly distributional attacks ($\kappa>1$) by supporting adversarial distributions on $2N$ points.
-This offers a complementary perspective to strong baselines such as AutoAttack and the RobustBench leaderboard (Croce & Hein, 2020; Croce et al., 2021).
-Empirically, WDA with $\kappa=2$ consistently finds stronger adversarial examples than state-of-the-art methods across diverse settings: achieving lower robust accuracy than APGD-DLR on CIFAR-10/100 with WideResNet backbone (Zagoruyko & Komodakis, 2016) on both $\ell_{\infty}$ and $\ell_{2}$ perturbations.
-When integrated into the Adaptive Auto Attack framework, our method matches or exceeds the ensemble performance of A3.
-These results demonstrate that the distributional perspective not only provides tighter theoretical certificates but also yields more effective attacks, validating our claim that existing robustness evaluations underestimate vulnerability by restricting to $\Omega_{\infty}$ rather than the larger $\Omega_{1}$ ambiguity set assumed by certificates.
+    Finally, we bridge the gap between WDRO theory and adversarial evaluation by introducing the Wasserstein Distributional Attack (WDA) and its adaptive variant (WDA++), which directly constructs adversarial distributions within the Wasserstein ball rather than restricting to point-wise perturbations.
+Unlike existing attacks that place all adversarial examples on the $\epsilon$-ball boundary, WDA flexibly interpolates between point-wise ($\kappa=1$) and truly distributional attacks ($\kappa>1$) by supporting adversarial distributions on $2N$ points, see Figure 2.
+WDA++ further optimizes the $2N$-support distribution via adaptive transport budget allocation.
+Empirically, WDA and WDA++ consistently find stronger adversarial examples than state-of-the-art methods across diverse settings (Table 1).
+When integrated into the Adaptive Auto Attack framework (Liu et al., 2022), our method matches or exceeds the ensemble performance of A3.
+Our results demonstrate that the distributional perspective not only provides tighter theoretical certificates but also yields more effective attacks.
+Crucially, this exposes the limitations of current evaluation standards (e.g., AutoAttack (Croce and Hein, 2020), RobustBench (Croce et al., 2021)) which significantly overestimate robustness by restricting to $\Omega_{\infty}$ rather than the larger $\Omega_{1}$ ambiguity set assumed by certificates.
+Figure 1: WDRO bounds and PGD attack loss for a ReLU classifier with $n=K=2$ and one hidden layer of dimension 8.
+Lower bound curves are the cumulative $\bm{l}$ as more reachable activation masks are considered.
 Figure 2: Left: Wasserstein ambiguity ball $\Omega_{p}=\left\{\mathbb{P}\colon\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\right\}$ inclusion and its admissible attacks.
@@
 2 Preliminaries
-Notations
+Notations.
 We denote basis vector as $\bm{e}_{k}$; indicator function as $\bm{1}_{\{\cdot\}}$; Dirac measure as $\bm{\delta}_{z}$; input dimension $n$, and output dimension as $K$.
 An empirical dataset is denoted $\{Z^{(1)},\dots,Z^{(N)}\}$ with $Z=(x,y)\in\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{K}$; empirical distribution $\mathbb{P}_{N}=\sum_{i}\mu_{i}\bm{\delta}_{Z^{(i)}}$ with $Z^{(i)}=(x^{(i)},y^{(i)}=\bm{e}_{k_{i}})$.
-Norms $\|\cdot\|_{r}$ and $\|\cdot\|_{s}$ are dual with 1/r+1/s=11/{r}+1/{s}=1.
+Norms $\|\cdot\|_{r}$ and $\|\cdot\|_{s}$ are dual with $1/r+1/s=1$.
 For a matrix $A$, $\|A\|_{{r}\to{s}}=\sup_{\|u\|_{r}=1}\|Au\|_{s}$.
-Rectifier $\cdot]_{+}=\max{\{0,\cdot\}}$.
+Rectifier $(\cdot)_{+}=\max{\{0,\cdot\}}$.
 Recession cone $\mathrm{rec}(\cdot)$.
@@
 Cross-entropy loss $\ell(x,y;\theta)=-\sum_{k=1}^{K}y_{k}\log\operatorname{softmax}(\theta(x))_{k}$.
-Analogous DLR loss as defined in Croce & Hein (2020).
-We define the dual-norm maximizer $\mathcal{M}_{r}$ by
+Analogous DLR loss as defined in Croce and Hein (2020).
+Dual-norm maximizer $\mathcal{M}_{r}$
 $\mathcal{M}_{r}\colon g\mapsto\arg\max_{h}\left\{\langle g,h\rangle\mid\|h\|_{r}=1\right\}=\left\{\begin{array}[]{ll}\operatorname{sign}(g)&\text{if }r=\infty,\\ g/\|g\|_{2}&\text{if }r=2,\\ \operatorname{sign}({g}_{k^{\prime}})\bm{e}_{k^{\prime}}\text{ with }k^{\prime}\in\arg\max_{k}|g_{k}|&\text{if }r=1,\end{array}\right$.
 (1)
-and projection $\Pi_{r,{x},\kappa\epsilon}$ by
-$\Pi_{r,{x},R}\colon x\mapsto\arg\min_{\xi}\left\{\|\xi-x\|_{2}^{2}\mid\|\xi\|_{r}\leq R\right\}.$ (2)
-Wasserstein Distributionally Robust Optimization (WDRO)
-Robustness guarantees and certificates aim to make model predictions trustworthy under adversarial manipulation (Wong & Kolter, 2018; Cohen et al., 2019; Salman et al., 2019).
-Empirical risk minimization model $\inf_{\theta}\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]$ optimizes average performance on the observed data but offers no protection against worst-case shifts nearby.
-Distributionally robust optimization (DRO) addresses this by choosing parameters that perform well against all distributions within a prescribed neighborhood: $\inf_{\theta}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)].$ Here, the worst-case loss is taken over all admissible distributions $\mathbb{P}\in\mathcal{P}$.
+and projector $\Pi_{r,{x},R}$
+[…]
+WDRO.
+Robustness certificates aim to make model predictions trustworthy under adversarial manipulation (Wong and Kolter, 2018; Cohen et al., 2019; Salman et al., 2019).
+The empirical risk minimization model $\inf_{\theta}\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]$ optimizes average performance on the observed data but offers no protection against worst-case shifts nearby.
+Distributionally robust optimization (DRO) addresses this by choosing parameters that perform well against all distributions within a prescribed neighbourhood: $\inf_{\theta}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)].$ Here, the worst-case loss is taken over all admissible distributions $\mathbb{P}\in\mathcal{P}$.
 The ambiguity (or uncertainty) set $\mathcal{P}$ is often constructed by collecting all distributions $\mathbb{P}$ that are similar to the empirical distribution $\mathbb{P}_{N}$.
-In this work, we focus on the Wasserstein ambiguity set, which is a ball centered at $\mathbb{P}_{N}$ under the Wasserstein distance.
-Given a ground cost $d$ on the space of data $\mathcal{Z}$, the Wasserstein distance (Villani, 2008) between two distributions $\mathbb{P}$ and $\mathbb{Q}$ it is defined as $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\left(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{Z}\times\mathcal{Z}}d^{p}(z^{\prime},z)\,\mathrm{d}\pi(z^{\prime},z)\right)^{1/p}\text{ for}\,p\in[1,\infty);$, and $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\operatorname{ess.sup}_{{\pi}}(d)\text{ for}\,p=\infty$.
+In this work, we focus on the Wasserstein ambiguity set, which is a ball centred at $\mathbb{P}_{N}$ under the Wasserstein distance.
+Given a ground cost $d$ on the space of data $\mathcal{Z}$, the Wasserstein distance $\mathcal{W}_{d,p}$ (Villani and others, 2008) between two distributions $\mathbb{P}$ and $\mathbb{Q}$ is defined for $p\in[1,\infty)$ as
+$\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\left(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{Z}\times\mathcal{Z}}d^{p}(z^{\prime},z)\,\mathrm{d}\pi(z^{\prime},z)\right)^{1/p}$
+and $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\operatorname{ess.sup}_{{\pi}}(d)$ for $p=\infty$.
 Intuitively, the Wasserstein distance between two distributions $\mathbb{P}$ and $\mathbb{Q}$ is defined as the minimum cost to transport the mass of $\mathbb{P}$ to $\mathbb{Q}$.
@@
 $\inf_{\theta}\sup_{\mathbb{P}\in\Omega_{p}}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\text{ where }\Omega_{p}=\left\{\mathbb{P}\mid\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\right\}.$ (3)
-It is worth noting that $\mathcal{W}_{d,p}\leq\mathcal{W}_{d,p^{\prime}}$ if $p\geq p^{\prime}$, thus $\Omega_{1}\supseteq\Omega_{p}\supseteq\Omega_{p^{\prime}}\supseteq\Omega_{\infty}$ (see Figure 1).
-Lipschitz Certificate
+It is worth noting that $\mathcal{W}_{d,p}\leq\mathcal{W}_{d,p^{\prime}}$ if $p\leq p^{\prime}$, thus $\Omega_{1}\supseteq\Omega_{p}\supseteq\Omega_{p^{\prime}}\supseteq\Omega_{\infty}$ (see Figure 2).
+For more details on Wasserstein distributionally robust optimization, we refer reader to Kuhn et al.
+(2019) and our Appendix A.
+Lipschitz Certificate.
 For $p=1$, the worst-case risk over a Wasserstein ball admits the standard Lipschitz upper bound
@@
 where $L$ is any Lipschitz constant of $z\mapsto\ell(z;\theta)$ with respect to the ground cost.
-This inequality follows from weak duality and is widely used to make the WDRO objective tractable: one replaces the inner maximization by the surrogate $L\epsilon$ and then controls $L$ (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024).
-In practice, estimating $L$ reduces to bounding the network’s (global or local) Lipschitz modulus, e.g., fast global products of per-layer operator norms (Virmaux & Scaman, 2018) or tighter activation-aware/local certificates (Jordan & Dimakis, 2020; Shi et al., 2022).
+This inequality follows from weak duality and is widely used to make the WDRO objective tractable: one replaces the inner maximization by the surrogate $L\epsilon$ and then controls $L$ (Mohajerin Esfahani and Kuhn, 2018; Blanchet et al., 2019; Gao and Kleywegt, 2023; Gao et al., 2024).
+In practice, estimating $L$ reduces to bounding the network’s (global or local) Lipschitz modulus, e.g., fast global products of per-layer operator norms (Virmaux and Scaman, 2018) or tighter activation-aware/local certificates (Jordan and Dimakis, 2020; Shi et al., 2022).
 Adversarial Attack.
-Adversarial attack methods often construct a perturbed distribution by shifting each sample $X^{(i)}$ along a specific adversarial direction $u^{(i)}$ to get $X_{\mathrm{adv}}^{(i)}$ (Goodfellow et al., 2014; Moosavi-Dezfooli et al., 2016; Carlini & Wagner, 2017).
-These methods are essentially point-wise attacks, which draws a distribution $\mathbb{P}_{\mathrm{adv}}=\sum_{i=1}^{N}\frac{1}{N}\bm{\delta}_{X^{(i)}_{\mathrm{adv}}}$ in the Wasserstein ambiguity set $\Omega_{p}=\{\mathbb{P}\colon\mathcal{W}_{d,p}\leq\epsilon\}$ when $p=\infty$ (see Figure 1).
+Adversarial attack methods often construct a perturbed distribution by shifting each sample $X^{(i)}$ along a specific adversarial direction $u^{(i)}$ to get $X_{\mathrm{adv}}^{(i)}$ (Goodfellow et al., 2014; Moosavi-Dezfooli et al., 2016; Carlini and Wagner, 2017).
+These methods are essentially point-wise attacks, which draws a distribution $\mathbb{P}_{\mathrm{adv}}={\scriptstyle\sum_{i=1}^{N}\frac{1}{N}}\bm{\delta}_{(X^{(i)}_{\mathrm{adv}},Y^{(i)})}$ in the Wasserstein ambiguity set $\Omega_{p}=\{\mathbb{P}\colon\mathcal{W}_{d,p}\leq\epsilon\}$ when $p=\infty$ (see Figure 2).
 Whereas, in the $p=1$ case, the ambiguity set only constrains the average transportation cost under an optimal coupling.
 Hence, the adversary may move some points farther and others less as long as the mean cost stays within budget.
-This creates a significant gap between the robustness measured against $\Omega_{p=\infty}$ attacks and the theoretical robustness or Lipschitz certificates 4 which are developed for $\Omega_{p=1}$ (Mohajerin Esfahani & Kuhn, 2018; Carlini et al., 2019; Rice et al., 2021).
+This creates a significant gap between the robustness measured against $\Omega_{p=\infty}$ attacks and the theoretical robustness or Lipschitz certificates equation 4 which are developed for $\Omega_{p=1}$ (Mohajerin Esfahani and Kuhn, 2018; Carlini et al., 2019; Rice et al., 2021).
 3 Tractable Interpretation of WDRO for Neural Networks
-For certain shallow and convex models (e.g., linear regression, support vector machines, etc.), the tractable representation of the WDRO problem 3 is well-established in the literature (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024).
+For certain shallow and convex models (e.g., linear regression, support vector machines, etc.), the tractable representation of the WDRO problem equation 3 is well-established in the literature (Mohajerin Esfahani and Kuhn, 2018; Blanchet et al., 2019; Gao and Kleywegt, 2023; Gao et al., 2024).
 This tractable form enables a computational advantage and provides a clear interpretation of the robustness of regularization mechanisms.
@@
 Remark.
-Consider a single-point empirical $\mathbb{P}_{N}=\bm{\delta}_{\{X^{(1)}=2\}}$ and a loss given by
+Consider a single-point empirical distribution $\mathbb{P}_{N}=\bm{\delta}_{\{X^{(1)}=2\}}$ and a loss function given by
 $\ell(x)=\begin{cases}|x|\quad\text{if }|x|\leq 1,\\ \frac{1}{2}|x|+\frac{1}{2}\quad\text{otherwise}.\end{cases}$
-Then $\ell$ is Lipschitz with modulus 1, however $\sup_{\Omega_{1}}\mathbb{E}_{\mathbb{P}}[\ell(X)]=\mathbb{E}_{\mathbb{P}_{N}}[\ell(X)]+\frac{\epsilon}{2}\quad\text{for any}\,\epsilon>0.$
-As presented in the following sections, our main theoretical results (Theorem 3.1 and 3.3) show that Lipschitz modulus provides a tight upper bound for the WDRO problem 3 for a class of ReLU neural networks and smooth activated neural networks.
+Then $\ell$ is Lipschitz with Lipschitz constant 1, however for any $\epsilon>0$,
+$\sup_{\Omega_{1}}\mathbb{E}_{\mathbb{P}}[\ell(X)]=\mathbb{E}_{\mathbb{P}_{N}}[\ell(X)]+\frac{1}{2}\epsilon.$
+In the following, our main theoretical results (Theorems 3.1 and 3.3) show that Lipschitz modulus provides a tight upper bound for the WDRO problem equation 3 for a class of ReLU neural networks and smooth activated neural networks.
 3.1 Exact and tractable interpretation of WDRO for ReLU neural networks
@@
 $L_{\theta}=\sup_{x\in\mathcal{X}}\sup_{J\in\partial\theta(x)}\|J\|_{r\rightarrow\tilde{r}}$, (5)
-where $J\in\partial\theta(x)$ is a general Jacobian of $\theta$ at $x$, then Jordan & Dimakis (2020, Theorem 1) has shown that $\|\theta(x^{\prime})-\theta(x)\|_{\tilde{r}}\leq L_{\theta}\times\|x^{\prime}-x\|_{r}$ for any $x^{\prime},x\in\mathcal{X}$.
-Moreover, if $\theta$ is in general position (Jordan & Dimakis, 2020, Definition 4), then the chain rule applies and any general Jacobian $J$ must has a form as $W_{H+1}D_{H}W_{H}\cdots D_{1}W_{1}$ for some $[0,1]$-diagonal matrix $D_{h}$, $h=1,\dots,H$.
-It is worth noting that the set of ReLU networks not in general position is negligible (Jordan & Dimakis, 2020, Theorem 3).
-Now in equation 5, the maximizer of a convex function (norm operator) must happen at vertices, thus we only need to consider $0/1$-diagonal matrix $D_{h}$.
+where $J\in\partial\theta(x)$ is a general Jacobian of $\theta$ at $x$, then Jordan and Dimakis (2020, Theorem 1) has shown the inequality that $\|\theta(x^{\prime})-\theta(x)\|_{\tilde{r}}\leq L_{\theta}\times\|x^{\prime}-x\|_{r}$ for any $x^{\prime},x\in\mathcal{X}$.
+Moreover, if $\theta$ is in general position (Jordan and Dimakis, 2020, Definition 4), then the chain rule applies and any general Jacobian $J$ must have a form as $W_{H+1}D_{H}W_{H}\cdots D_{1}W_{1}$ for some $[0,1]$-diagonal matrix $D_{h}$, $h=1,\dots,H$.
+It is worth noting that the set of ReLU networks not in general position is negligible (Jordan and Dimakis, 2020, Theorem 3).
+Now in equation 5, the maximization of a convex function (norm operator) is achieved at the vertices, thus we only need to consider $0/1$-diagonal matrix $D_{h}$.
 We formally introduce the concept of mask as follows.
@@
 $\operatorname{pre}_{h}\colon x\mapsto W_{h}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{h}).$
-Given this definition of mask and note that $\mathcal{D}_{\mathcal{X}}$ is finite, one can rewrite equation 5 as $L_{\theta}=\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\|J_{\bm{D}}\|_{r\rightarrow\tilde{r}}$.
-We adopt this notion and show that it induces an upper bound for the Wasserstein distributional robust optimization (WDRO) problem 3 with cross-entropy loss.
+Given this definition, and noting that $\mathcal{D}_{\mathcal{X}}$ is finite, one can rewrite equation 5 as $L_{\theta}=\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\|J_{\bm{D}}\|_{r\rightarrow\tilde{r}}$.
+We adopt this notion and show that it induces an upper bound for the Wasserstein distributional robust optimization (WDRO) problem equation 3 with cross-entropy loss.
 Moreover, this upper bound is tight for a class of monotonic ReLU networks.
-Theorem 3.1 (WDRO for ReLU).
-Given a ReLU network $\theta(x)=W_{H+1}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{H})$ being in general position, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
-$\bm{L}\triangleq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}},$ (6)
+Theorem 3.1 (WDRO for ReLU Networks).
+Given a ReLU network
+$\theta(x)=W_{H+1}(\operatorname{ReLU}(\cdots(W_{1}x+b_{1})\cdots)+b_{H})$
+being in general position, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
+$\bm{L}\triangleq 2^{1/r}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}},$ (6)
 and
@@
 $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon\leq\sup_{\mathbb{P}\colon\mathcal{W}_{d,1}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\leq\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{L}\epsilon.$ (8)
-Moreover, if the dual-norm maximizer $\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of 7, and $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})$ is the largest increment direction of $J_{\bm{D}^{\star}}$, then $\bm{l}=\bm{L}$.
+Moreover, let $\bm{D}^{\star}$ be a maximizer of equation 6, $(k^{\prime\star},k^{\star})$ be a maximizer of equation 7 and $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})$ be the largest increment direction of $J_{\bm{D}^{\star}}$.
+If the dual-norm maximizer $\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ then $\bm{l}=\bm{L}$.
 Proof.
-To prove inequality 8, we show that $\ell(\cdot,\theta)$ is $\bm{L}$-Lipschitz, and a direction $u$ found in equation 7 induces an admissible attack $\mathbb{P}_{\mathrm{adv}}$ satisfying that $\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$ and $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$.
+To prove inequality equation 8, we show that the loss function $\ell(\cdot,\theta)$ is $\bm{L}$-Lipschitz, and a direction $u$ found in equation 7 induces an admissible attack distribution $\mathbb{P}_{\mathrm{adv}}$ satisfying $\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$, $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$.
 To verify the sufficient condition of $\bm{l}=\bm{L}$, we show that the constructed $\mathbb{P}_{\mathrm{adv}}$ provides $\bm{l}=\bm{L}$.
-We provide detailed proof in Appendix A.1.
+We provide detailed proof in Appendix B.1.
 ∎
@@
 While equation 7 provides a tight lower bound of the WDRO, it is impractical to scan through all $x\in\mathcal{X}$ and its mask $\mathcal{D}_{x}$.
-We then introduce a practical lower bound of which we consider the mask associated with the sample points only.
-Figure 2: WDRO bounds and PGD attack loss for a fixed $n=K=2$ ReLU classifier with one hidden layer of dimension 8.
-Lower-bound curves are the cumulative $\bm{l}$ as more reachable activation masks are considered.
+We then introduce a practical lower bound, of which we consider the mask associated with the sample points only.
 Corollary 3.2 (Practical lower bound).
 Given assumptions and notations used in Theorem 3.1, let $\mathcal{Z}_{N}=\{(X^{(1)},Y^{(1)}),\dots,(X^{(N)},Y^{(N)})\}$ and
-$\bm{l}_{N}\triangleq\max_{\begin{subarray}{c}(X^{(i)},Y^{(i)})\in\mathcal{Z}_{N},\\ \bm{D}\in\mathcal{D}_{x}\end{subarray}}\max_{k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k}-Y^{(i)})^{\top}J_{\bm{D}}u\mid u\in\mathrm{int}(\mathrm{rec}({\mathcal{C}_{\bm{D}}}))\right\}.$ (9)
+$\bm{l}_{N}\triangleq\max_{\begin{subarray}{c}(X^{(i)},Y^{(i)})\in\mathcal{Z}_{N}\\ \bm{D}\in\mathcal{D}_{x}\end{subarray}}\max_{k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k}-Y^{(i)})^{\top}J_{\bm{D}}u\mid u\in\mathrm{int}(\mathrm{rec}({\mathcal{C}_{\bm{D}}}))\right\}.$ (9)
 Then $\bm{l}_{N}\leq\bm{l}$.
-In the spirit of formulation 9, we propose a practical adversarial attack in Section 4 which aims to find adversarial direction $u^{(i)}$ for each sample $i$ so that it maximizes the change of the logit function.
+Based on the proof of our lower bound equation 17, we construct a worst-case distribution by moving mass from a sample along a direction $u$ that maximizes the margin term in equation 7.
+In Section 4, based on formulation equation 9, we create this construction empirically via the attack distribution equation 11 by choosing adversarial direction $u^{(i)}$ for each sample $i$ so that it maximizes the first-order increase of the corresponding logit margin.
 3.2 Exact and tractable interpretation of WDRO for smooth activation neural networks
-For networks with smooth activations, e.g, GELU (Hendrycks & Gimpel, 2016), SiLU/Swish (Ramachandran et al., 2017; Elfwing et al., 2018), WDRO duality connects worst-case (adversarial) risk to first-order geometry via the Jacobian of the logit map, yielding global Lipschitz-type upper penalties of the form $\sup_{x\in\mathcal{X}}\|J(x)\|_{r\to s}$.
-Compared to piecewise-linear ReLU certificates, smooth nets trade exact cell-wise constancy for differentiability along rays and curves, suggesting bounds driven by asymptotic Jacobian behavior rather than activation masks.
-Let $t$ be a positive scalar, $\theta:\mathcal{X}\subseteq\mathbb{R}^{n}\to\mathbb{R}^{K}$ be a classifier with smooth activations and cross-entropy loss; let $J(x)\in\mathbb{R}^{K\times n}$ be its Jacobian.
-We then have the following result.
+While ReLU networks require analyses of activation masks, modern architectures often employ smooth activation functions (e.g., GELU (Hendrycks and Gimpel, 2016), SiLU (Ramachandran et al., 2017)).
+In these settings, the WDRO duality connects the worst-case adversarial risk directly to the first-order geometry via the Jacobian of the logit map.
 Theorem 3.3 (WDRO for Smooth Networks).
-Let $\theta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$ be a differentiable network, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
-$\bm{L}\triangleq{2^{1/{s}}}\sup_{x\in{\mathcal{X}}}\left\|\nabla_{x}\theta(x)\right\|_{{r}\to{s}},$
+Given a differentiable network $\theta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{K}$ with $J_{\theta}(x)\in\mathbb{R}^{K\times n}$ being its Jacobian, $1/r+1/s=1$ and $\ell$ being the cross-entropy or DLR loss, define
+$\bm{L}\triangleq 2^{1/r}\sup_{x\in{\mathcal{X}}}\left\|J_{\theta}(x)\right\|_{{r}\to{s}},$
 and
-$\bm{l}\triangleq\sup_{x\in{\mathcal{X}}}\max_{k^{\prime}\neq k}\sup_{\|u\|_{r}=1}\left\{(\bm{e}_{k^{\prime}}-\bm{e}_{k})^{\top}\nabla_{x}\theta(x)u\right\}.$
+$\bm{l}\triangleq\sup_{x\in{\mathcal{X}}}\max_{k^{\prime}\neq k}\sup_{\|u\|_{r}=1}(\bm{e}_{k^{\prime}}-\bm{e}_{k})^{\top}J_{\theta}(x)u.$
 Then for any $\epsilon>0$,
 $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon\leq\sup_{\mathbb{P}\colon\mathcal{W}_{d,1}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\leq\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{L}\epsilon.$
-In this setting, first-order WDRO penalties are controlled by how $J(x)$ amplifies unit directions and how that amplification projects onto the most competitive non-true logit.
-The upper slope $\bm{L}$ is the global worst-case amplification, while the lower, margin–directional slope $\bm{l}$ follows rays $x^{(i)}+tu$ and harvests only the component along $(\bm{e}_{k^{\prime}}-\bm{e}_{k_{i}})$.
-When a ray both attains the global operator norm and aligns with a margin difference, the bound is tight to first order ($\bm{l}=\bm{L}$).
-This motivates the adversarial procedure used our adversarial attack algorithm WDA (Algogirthm 1), where we search for a direction $u$ and a rival class $k^{\prime}$ that maximize the first-order increase $(\bm{e}_{k^{\prime}}-\bm{e}_{k_{i}})^{\top}J(x)u$.
-4 Wasserstein Distributional Attack
-Existing point-wise attacks such as FGSM (Goodfellow et al., 2014), DeepFool (Moosavi-Dezfooli et al., 2016), AA (Croce & Hein, 2020), AAA (Liu et al., 2022), keep the adversarial distribution supported on exactly $N$ points, where each point $X^{(i)}_{\operatorname{adv}}$ is perturbed to be precisely on the boundary of the $\epsilon$-ball centered at $X^{(i)}$.
-To address this issue, we propose a novel method called Wasserstein Distributional Attack (WDA).
-At a high level, WDA constructs an adversarial distribution, $\mathbb{P}_{\mathrm{adv}}$, supported on a set of $2N$ points.
-This set comprises of $N$ original empirical samples $X^{(i)}$ and $N$ corresponding adversarial points $X^{(i)}_{\mathrm{adv}}$, each perturbed to an $r$-norm distance $\kappa\epsilon$ from $X^{(i)}$ using the first-order, margin-aligned directions predicted by Theorems 3.13.3, for some $\kappa\geq 1$.
-In other words,
-$\mathbb{P}_{\mathrm{adv}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa}\bm{\delta}_{(X_{\mathrm{adv}}^{(i)},Y^{(i)})}.$ (10)
-In the special case where $\kappa=1$, our proposed attack reduces to existing point-wise methods.
-When $\kappa=2$, WDA simplifies to a uniform distribution over all $2N$ points, with each point receiving a weight of $\frac{1}{2N}$.
-This $2N$-support mixture belongs to the $\Omega_{1}$ ambiguity set and serves as a constructive, distributional adversary; it is not necessarily the inner maximizer of WDRO.
-We now make the first-order ascent directions explicit; this is the step used by WDA to realize the margin-aligned rays from Theorems 3.13.3.
-Define (sub)gradient $g_{j}(x)\in\partial_{x}\big(\theta_{j}-\theta_{k}\big)(x)$.
-Then $\mathcal{M}_{r}(g_{j})$ give the per–iteration, first–order version of the ray ascent used in Theorems 3.13.3: within a ReLU cell (affine logits) or for smooth activations (continuous $J$), moving along $u_{j}=\mathcal{M}_{r}(g_{j})$ increases the gap at rate $\|g_{j}(x)\|_{s}$.
-During an initial probing phase, we evaluate all rivals $j\neq k$ using these first-order steps.
-At the end of that phase, we fix a single rival $j^{*}$ based on the logits magnitude and continue the remaining iterations.
-If we allow $j^{*}$ to change at every step, the update can oscillate across classes and chase locally steep but globally suboptimal directions for misclassifications.
-Finally, we project each step to the ball of radius $\kappa\varepsilon$ around the anchor $X^{(i)}$ to the WDRO budget.
-The procedure for implementing the Wasserstein Distributional Attack is presented in Algorithm 1.
-A visualization of our algorithm is shown in Figure 3.
-Algorithm 1 Wasserstein Distributional Attack (WDA)
-1:Inputs: neural network $\theta:\mathbb{R}^{n}\to\mathbb{R}^{K}$, empirical distribution $\mathbb{P}_{N}=\sum_{i=1}^{N}\bm{\delta}_{(X^{(i)},Y^{(i)})}$, budget $\epsilon>0$, cost-norm $r\in\{1,2,\infty\}$, WDA paramter $\kappa\geq 1$, step size $\alpha>0$, and $0<$ prob $\leq$ maxiter
-2:Outputs: Wasserstein distributional attack $\mathbb{P}_{\mathrm{adv}}$ such that $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$ where $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$
-3:Initialize: dual-norm maximizer $\mathcal{M}$ (1), projection $\Pi$ (2)
-4:for $i=1$ to NN do
-5:  $x_{0}\leftarrow X^{(i)}$, $\bm{e}_{k}\leftarrow Y^{(i)}$ for some $k=1,\dots,K$
-6:  for iter =0=0 to maxiter do
-7:   if iter << prob then $\mathcal{J}=\{1,\dots,K\}\setminus\{k\}$ else $\mathcal{J}=\{j^{*}\}$
-8:   $g_{j}\leftarrow\nabla_{x}\theta\left(x_{\mathrm{iter}}\right)^{\top}(\bm{e}_{j}-\bm{e}_{k})$ for $j\in\mathcal{J}$
-9:   $u_{j}\leftarrow\mathcal{M}_{r}(g_{j})$ for $j\in\mathcal{J}$
-10:   $\varphi_{j}\leftarrow\Pi_{r,X^{(i)},\kappa\epsilon}\left(x_{\mathrm{iter}}+\alpha u_{j}\right)$ for $j\in\mathcal{J}$
-11:   $j^{*}=\arg\max_{j\in\mathcal{J}}\theta_{j}\left(\varphi_{j}\right)$
-12:   $x_{\mathrm{iter+1}}\leftarrow\varphi_{j^{*}}$
-13:  end for
-14:  $X^{(i)}_{\mathrm{adv}}\leftarrow x_{\mathrm{maxiter}}$
-15:end for
-[…]
-17:return $\mathbb{P}_{\mathrm{adv}}$
-Figure 3: Wasserstein Distributional Attack (WDA, Alg.
-1) for $r=2$.
+In addition, this result extends certification guarantees to modern differentiable Transformer architectures such as GPT (Radford et al., 2019), ViT (Dosovitskiy et al., 2021), and Swin Transformer (Liu et al., 2021), which are often Pre-LN Transformers Xiong et al.
+(2020).
+Corollary 3.4 (WDRO for Pre-LN Transformers).
+Given assumptions and notations used in Theorem 3.3, if $\theta$ is a depth-$H$ (Pre-LN) Transformer, then by Castin et al. (2024, Theorem 3.3 and Lemma 3.8), the upper bound satisfies
+$\bm{L}=2^{1/r}\sup_{x\in\mathcal{X}}\|J_{\theta}(x)\|_{r\to s}\leq 2^{1/r}(nK)^{(\frac{1}{2}-\frac{1}{r})_{+}}\times\Lambda_{\theta},$ (10)
+where $\Lambda_{\theta}\triangleq L_{\rm head}L_{\rm emb}\prod_{h=1}^{H}\left(1+L_{\rm LN}^{(h)}L_{\rm MHA}^{(h)}\right)\left(1+L_{\rm LN^{\prime}}^{(h)}L_{\rm FFN}^{(h)}\right)$, and all module constants are Euclidean Lipschitz constants on $\mathcal{X}$.
+Moreover, their explicit bound of the form $L_{\rm MHA}^{(h)}=\mathcal{O}(\sqrt{n})$ implies $\bm{L}=\mathcal{O}\left(2^{1/r}(nK)^{(\frac{1}{2}-\frac{1}{r})_{+}}\times n^{H/2}\right)$ up to architecture and parameter-dependent constants.
+4 Wasserstein Distributional Attacks
+Existing point-wise attacks such as FGSM (Goodfellow et al., 2014), DeepFool (Moosavi-Dezfooli et al., 2016), AA (Croce and Hein, 2020), AAA (Liu et al., 2022), keep the adversarial distribution supported on exactly $N$ points, where each point $X^{(i)}_{\operatorname{adv}}$ is perturbed to be precisely on the boundary of the $\epsilon$-ball centred at $X^{(i)}$.
+To address this issue, we propose a novel method called Wasserstein Distributional Attacks which constructs an adversarial distribution $\mathbb{P}_{\mathrm{adv}}$, supported on a set of $2N$ label-preserving points
+$\mathbb{P}_{\mathrm{adv}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa_{i}}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa_{i}}\bm{\delta}_{(X^{(i)}_{\mathrm{adv}},Y^{(i)})},$ (11)
+where $\kappa_{i}\geq 1$.
+(See Figure 2.) This $2N$-support consists of $N$ original empirical samples $X^{(i)}$ and $N$ corresponding adversarial points $X^{(i)}_{\mathrm{adv}}$, each perturbed using the first-order, margin-aligned directions predicted by Theorems 3.1 and 3.3.
+WDA.
+WDA corresponds to $\kappa_{i}\equiv\kappa$.
+In the special case where $\kappa=1$, the attack reduces to point-wise methods, while $\kappa=2$ yields a uniform distribution over all $2N$ points, with each point receiving a weight of $\frac{1}{2N}$.
+The mixture equation 11 lies in $\Omega_{p}$ whenever $\|X^{(i)}_{\mathrm{adv}}-X^{(i)}\|_{r}\leq\kappa^{1/p}\epsilon$ for all $i$ (Corollary C.2) and serves as a constructive and feasible distributional adversary.
+WDA++.
+WDA++ keeps the same two-point anchored form equation 11 but chooses the mixing weight adaptively.
+For each sample $i$ we calculate $d_{i}$, an estimate of the closest distance to flip an image’s label and set $1/\kappa_{i}=\min\{1,t_{i}/d_{i}^{p}\}$ for a transport allocation $t_{i}\geq 0$ (with the conventions $1/\kappa_{i}=1$ if $d_{i}=0$ and $1/\kappa_{i}=0$ if $d_{i}=\infty$), yielding $X^{(i)}_{\rm adv}$ chosen near the closest flip.
+By Corollary C.2, $\mathcal{W}_{d,p}(\mathbb{P}_{\rm adv},\mathbb{P}_{N})^{p}\leq\sum_{i}\mu_{i}t_{i}$, hence $\mathbb{P}_{\rm adv}\in\Omega_{p}$ whenever $\sum_{i}\mu_{i}t_{i}\leq\epsilon^{p}$.
+Full implementation details are deferred to Appendix E.
+Figure 3: Wasserstein Distributional Attack (WDA, Algorithm 1) for $r=2$.
 At each iteration $x_{t}$, WDA forms K1K\!-\!1 candidates $\varphi_{j}$ and updates using the one with the largest logit $\theta_{j}(\varphi_{j})$.
 For reference, PGD follows the dual-norm gradient direction; DeepFool linearizes the decision boundary.
+Algorithm 1 Wasserstein Distributional Attack (WDA)
+Inputs: neural network $\theta:\mathbb{R}^{n}\to\mathbb{R}^{K}$, empirical distribution $\mathbb{P}_{N}=\frac{1}{N}\sum_{i=1}^{N}\bm{\delta}_{(X^{(i)},Y^{(i)})}$, budget $\epsilon>0$, cost-norm $r\in\{1,2,\infty\}$, Wasserstein order $p\in\{1,2\}$, WDA parameter $\kappa\geq 1$, step size $\alpha>0$, and $0<$ probe $\leq$ maxiter
+Outputs: Wasserstein distributional attack $\mathbb{P}_{\mathrm{adv}}$ such that $\mathcal{W}_{d,p}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$ where $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$
+Initialize: dual-norm maximizer $\mathcal{M}$ equation 1, projection $\Pi$ equation 2
+for $i=1$ to NN do
+  $x_{0}\leftarrow X^{(i)}$, $\bm{e}_{k}\leftarrow Y^{(i)}$ for some $k=1,\dots,K$
+  for iter =0=0 to maxiter do
+   if iter << probe then $\mathcal{J}=\{1,\dots,K\}\setminus\{k\}$ else $\mathcal{J}=\{j^{*}\}$
+   $g_{j}\leftarrow\nabla_{x}\theta\left(x_{\mathrm{iter}}\right)^{\top}(\bm{e}_{j}-\bm{e}_{k})$ for $j\in\mathcal{J}$
+   $u_{j}\leftarrow\mathcal{M}_{r}(g_{j})$ for $j\in\mathcal{J}$
+   $\varphi_{j}\leftarrow\Pi_{r,X^{(i)},\kappa^{1/p}\epsilon}\left(x_{\mathrm{iter}}+\alpha u_{j}\right)$ for $j\in\mathcal{J}$
+   $j^{*}\leftarrow\arg\max_{j\in\mathcal{J}_{i}}\left(\theta_{j}(\varphi_{j})-\theta_{k}(\varphi_{j})\right)$
+   $x_{\mathrm{iter+1}}\leftarrow\varphi_{j^{*}}$
+  end for
+  $X^{(i)}_{\mathrm{adv}}\leftarrow x_{\mathrm{\texttt{maxiter}}}$
+end for
+$\mathbb{P}_{\mathrm{adv}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa}\bm{\delta}_{(X_{\mathrm{adv}}^{(i)},Y^{(i)})}$
+return $\mathbb{P}_{\mathrm{adv}}$
+Adversarial direction.
+Define the (sub)gradient $g_{j}(x)\in\partial_{x}\big(\theta_{j}-\theta_{k}\big)(x)$.
+Then $\mathcal{M}_{r}(g_{j})$ provides the per-iteration, first-order instance of the ray ascent from Theorems 3.1 and 3.3: within a ReLU cell or for smooth activations, moving along $u_{j}=\mathcal{M}_{r}(g_{j})$ increases the gap at rate $\|g_{j}(x)\|_{s}$.
+During an initial probing phase, we evaluate all rivals $j\neq k$ using these first-order steps.
+At the end of this phase, we fix a single rival $j^{*}$ based on the logit magnitude and continue the remaining iterations.
+Fixating $j^{*}$ prevents the update from oscillating across classes or chasing locally steep but globally suboptimal directions.
+Finally, we project each step to the ball of radius $\kappa^{1/p}\varepsilon$ around the anchor $X^{(i)}$.
+A visualization of our algorithm is shown in Figure 3.
+The procedures for implementing WDA and WDA++ are presented in Algorithms 1 and 2.
 5 Related Work
-Robustness Certificates
-Early scalable global certificates control the Lipschitz constant by multiplying per-layer operator norms, which is fast to compute yet data-agnostic and typically loose on deep nets (Virmaux & Scaman, 2018).
+Robustness certificates.
+Early scalable global certificates control the Lipschitz constant by multiplying per-layer operator norms, which is fast to compute yet data-agnostic and typically loose on deep nets (Virmaux and Scaman, 2018).
 For ReLU networks, local (activation-aware) methods exploit piecewise linearity to produce much tighter, input-conditioned certificates on individual activation regions (Katz et al., 2017; Ehlers, 2017; Weng et al., 2018; Singh et al., 2018; Shi et al., 2022).
-Most relevant to exact local Lipschitzness, Jordan & Dimakis (2020) showed that for a broad class of ReLU networks in general position, the local Lipschitz constant can be computed exactly by optimizing over activation patterns.
-Adversarial Attacks
+Most relevant to exact local Lipschitzness, Jordan and Dimakis (2020) showed that for a broad class of ReLU networks in general position, the local Lipschitz constant can be computed exactly by optimizing over activation patterns.
+Adversarial attacks.
 Adversarial Attack methods seek for perturbation $x^{\prime}$ formed by adding a small, human-imperceptible perturbation to a clean input $x$ that causes misclassification (Szegedy et al., 2014).
@@
 black-box), the admissible perturbation set (e.g., $r_{2}$ balls with budget $\epsilon$), and the objective (e.g., worst-case loss within the ball).
-Canonical white-box methods include FGSM (Goodfellow et al., 2014), multi-step PGD (Madry et al., 2018), CW (Carlini & Wagner, 2017), and gradient-based margin attacks such as DeepFool (Moosavi-Dezfooli et al., 2016).
+Canonical white-box methods include FGSM (Goodfellow et al., 2014), multistep PGD (Madry et al., 2018), CW (Carlini and Wagner, 2017), and gradient-based margin attacks such as DeepFool (Moosavi-Dezfooli et al., 2016).
+In the context of distributional threats, Bai et al.
+(2023) introduced W-PGD which transports the empirical measure while maintaining its support size.
+This effectively restricts the attack to a point-wise one.
+Our work adopts a different perspective by expanding the support to $2N$ points, thereby exploring a larger region of the ambiguity set.
 Decision-based and score-free attacks (black-box) include Boundary Attack (Brendel et al., 2021) and Square Attack (Andriushchenko et al., 2020).
 Robust evaluation is subtle: gradient masking can inflate apparent robustness if attacks are not adapted (Athalye et al., 2018).
-To standardize evaluation, AutoAttack (AA) (Croce & Hein, 2020) composes strong, parameter-free attacks (APGD-CE, APGD-DLR, FAB, Square) and is widely adopted for reporting robust accuracy.
+To standardize evaluation, AutoAttack (AA) (Croce and Hein, 2020) composes strong, parameter-free attacks (APGD-CE, APGD-DLR, FAB, Square) and is widely adopted for reporting robust accuracy.
 RobustBench (Croce et al., 2021) curates model zoos and standardized test protocols across datasets and $r_{p}$ threat models, enabling comparable and reproducible robustness claims.
 Liu et al.
-(2022) proposed Adaptive Auto Attack (A3), which incorporates Adaptive Direction Initialization (ADI) and Online Statistics-based Discarding (ODS) (Tashiro et al., 2020) to enhance attack efficiency.
+(2022) proposed Adaptive Auto Attack (A3), which incorporates Adaptive Direction Initialization (ADI) and Online Statistics-based Discarding (ODS) (Tashiro et al., 2020) with a PGD framework to enhance attack efficiency.
 In our experiments, we report robustness under AA and A3 following RobustBench conventions and use them as baselines for comparison.
 Several works have focused on adversarial attacks tailored to ReLU networks.
-Croce & Hein (2018) introduced rLR-QP, a gradient-free method that navigates the piecewise-linear regions of ReLU models by solving convex subproblems and enhancing exploration with randomization and local search.
+Croce and Hein (2018) introduced rLR-QP, a gradient-free method that navigates the piecewise-linear regions of ReLU models by solving convex subproblems and enhancing exploration with randomization and local search.
 More recently, Zhang et al.
 (2022) developed BaB-Attack, a branch-and-bound framework that operates in activation space, leveraging bound propagation, beam search, and large neighborhood search to uncover stronger adversarial examples than conventional gradient-based approaches, particularly on hard-to-attack inputs.
-6 Experiments
-Experimental settings
-To evaluate the effectiveness of WDA, we test the adversarial robustness of several state-of-the-art defense models on the CIFAR-10 dataset.
-For comparison, we report both $r_{\infty}$ and $r_{2}$ robustness under perturbation budgets of $\epsilon=8/255$ and $\epsilon=0.5$, respectively.
-In addition to point-wise attack, we conduct a separate Wasserstein distributional attack experiment to further validate our method.
-Specifically, we set $\kappa=2$ in Algorithm 1 to find the adversarial (distributional) attack $\mathbb{P}_{\mathrm{adv}}$equation 10 and reporting classification accuracy on the distribution by $(1-1/\kappa)\times\mathrm{cleanaccuracy}+\kappa\times\mathrm{adversarialaccuracy}$.
-Our attack is benchmarked against AA, APGD-DLR, APGD-CE (Croce & Hein, 2020), and A3 (Liu et al., 2022).
-The defense models, along with their official implementations and pretrained weights, are obtained from RobustBench Croce et al.
-(2021).
-All experiments are conducted on 2x NVIDIA GeForce RTX 4090 GPU and 1x NVIDIA H200.
-6.1 Comparison with existing baselines
-Setup
-For the baselines AA, APGD-DLR, APGD-CE, and A3, we adopt the configurations reported in their respective research papers.
-Meanwhile, in WDA, we set the number of probe steps to $\alpha_{probe}=10$, and use $\alpha_{atk}=20$ attack iterations.
-We further propose A3++, an extension of A3 that incorporates our attack into its framework.
-Table 1: Comparison of robust accuracy of WDA and baseline methods against various defenses on CIFAR-10 with $r_{\infty}$ perturbations ($\epsilon=8/255$) and $r_{2}$ perturbations ($\epsilon=0.5$), as well as on CIFAR-100 with $r_{\infty}$ perturbations ($\epsilon=8/255$).
-The first column reports the source paper corresponding to each defense method.
-Paper Model Clean Single method Ensemble method
-APGD -CE APGD -DLR WDA ($\kappa=1$) WDA ($\kappa=2$) AA A3 A3++
+As pointed out in Zhang et al.
+(2022); Croce et al.
+(2020), rLR-QP and BaB-Attack are not as efficient as gradient based attack, therefore, for single method ($\Omega_{\infty}$) baseline we only use APGD.
+6 Comparison with Existing Baselines
+Experimental setup.
+We evaluate the effectiveness of WDA and WDA++ in amplifying misclassification rates against state-of-the-art defences on CIFAR-10, CIFAR-100, and ImageNet.
+We report robustness under $r_{\infty}$ and $r_{2}$ perturbations with budgets $\epsilon\in\{4/255,8/255,0.5\}$.
+Defence models and pre-trained weights are sourced from RobustBench (Croce et al., 2021).
+Our benchmarks include single-method baselines (APGD-CE, APGD-DLR, W-PGD) and ensemble methods (AA, A3).
+For baselines, we copy the official default configurations.
+For WDA, we use probe steps $\texttt{probe}=10$, attack iterations $\texttt{maxiter}=20$ and step size $\alpha=0.2$ for $r_{2},\epsilon=0.5$, $\alpha\in\{0.01,0.02\}$ for $r_{\infty},\epsilon\in\{4/255,8/255\}$ respectively.
+We further propose A3-WDA, an extension that substitutes the PGD attack of A3 with our WDA attack.
+In addition to standard point-wise evaluation with $\kappa=1$, we validate the distributional threat model using $\kappa=2$.
+We report the classification accuracy on the constructed distribution $\mathbb{P}_{\mathrm{adv}}$ as $(1-1/\kappa)\times\mathrm{acc_{clean}}+(1/\kappa)\times\mathrm{acc_{adv}}$.
+Comprehensive ablation studies are detailed in Appendix D.
+For WDA++, we evaluate under the constrained $\Omega_{2}$ set (comparable to W-PGD) and the full Wasserstein ambiguity ball $\Omega_{1}$.
+We adopt the parameter settings of WDA and detail the implementation in Appendix E.
+All experiments are conducted on 2x NVIDIA GeForce RTX 4090 and 1x NVIDIA H200 GPUs.
+Results on CIFAR-10.
+Table 1 presents performance on CIFAR-10.
+Under both $r_{\infty}$ and $r_{2}$ threat models, WDA++ ($\Omega_{1}$) significantly reduces robust accuracy compared to all other methods.
+This confirms that standard robustness benchmarks significantly overestimate true robustness against distributional shifts.
+Moreover, expanding the ambiguity set to the full Wasserstein ball ($\Omega_{1}$) exposes vulnerabilities that both distributional adversaries under $\Omega_{2}$ and point-wise attacks under $\Omega_{\infty}$ fail to detect.
+Within the $\Omega_{2}$ constraint, WDA++ produces lower robust accuracy compared to W-PGD, confirming that our attack finds better directions in the loss landscape and utilises the Wasserstein budget $\epsilon$ more efficiently.
+Within the point-wise setting ($\kappa=1$), WDA remains competitive, outperforming single-method attacks (APGD-CE and APGD-DLR) and achieves results comparable to ensemble-based methods, indicating its ability to match the strength of more computationally demanding attack aggregations.
+When integrated into the ensemble framework (A3-WDA), our method consistently achieves the lowest robust accuracy, surpassing the standard A3 and AA benchmarks.
+Table 1: Comparison of robust accuracy of WDA and baseline methods against various defences on CIFAR-10, CIFAR-100 and ImageNet.
+The best (lowest) robust accuracy of single methods and ensemble methods are highlighted in underline and bold, respectively.
 […]
-Bartoldson et al. (2024) WRN-94-16 93.68 76.15 74.31 74.05 65.25 73.71 73.55 73.54
-Bartoldson et al. (2024) WRN-82-8 93.11 74.17 72.54 71.85 62.06 71.59 71.46 71.46
-Cui et al. (2024) WRN-28-10 92.16 70.60 68.62 68.07 60.01 67.73 67.58 67.57
-Wang et al. (2023) WRN-70-16 93.25 73.46 71.68 71.02 63.08 70.69 70.53 70.52
-Wang et al. (2023) WRN-28-10 92.44 70.24 68.24 67.60 60.96 67.31 67.17 67.17
-Xu et al. (2023) WRN-28-10 93.69 67.08 69.00 66.39 63.25 63.89 63.93 63.84
-Sehwag et al. (2022) RN-18 84.59 58.40 57.66 56.30 54.65 55.54 55.50 55.50
-[…]
-Wang et al. (2023) WRN-70-16 95.54 85.66 85.30 85.00 77.63 84.97 84.96 84.97
-Wang et al. (2023) WRN-28-10 95.16 84.52 83.89 83.71 76.31 83.68 83.68 83.68
-Sehwag et al. (2022) WRN-34-10 90.93 78.23 78.16 77.51 72.01 77.24 77.22 77.25
-Sehwag et al. (2022) RN-18 89.76 75.24 75.32 74.69 69.75 74.41 74.41 74.40
-Ding et al. (2020) WRN-28-4 88.02 66.62 66.62 66.22 63.04 66.09 66.05 66.06
-Cui et al. (2024) WRN-28-10 89.05 66.58 67.08 66.59 64.19 66.44 66.41 66.42
-[…]
-Wang et al. (2023) WRN-28-10 72.58 44.09 39.66 39.12 43.61 38.77 38.70 38.71
-Addepalli et al. (2022) RN-18 65.45 33.47 28.82 28.26 37.64 27.67 27.65 27.63
-Cui et al. (2024) WRN-28-10 73.85 43.82 40.37 39.57 43.68 39.18 39.17 39.14
-Robustness on $r_{\infty}$ and $l_{2}$ Perturbations for CIFAR-10
-Table 1 presents the robust accuracy of several attack methods under $r_{\infty}$ perturbations with $\epsilon=8/255$ and $r_{2}$ perturbations with $\epsilon=0.5$.
-Across both single and ensemble based threat models, WDA consistently outperforms other single-method attacks (APGD-CE and APGD-DLR), highlighting its effectiveness as a stronger standalone adversarial evaluation.
-Moreover, WDA achieves results that are often comparable to ensemble-based methods, indicating its ability to match the strength of more computationally demanding attack aggregations.
-Within the ensemble family, A3++ demonstrates clear improvements over AA and provides competitive performance with A3, surpassing it on several defense models (3 out of 7 under $r_{\infty}$ and 1 out of 6 under $r_{2}$).
-Notably, under $r_{\infty}$ and $r_{2}$, WDA ($\kappa=2$) produces lower robust accuracy values than other attacking methods across all defense models.
-This highlights the potential of Wasserstein distributional attack.
-6.2 Ablation study
-Figure 4: Robust accuracy with varying $\kappa$ on different defense methods.
-Figure 5: Convergence of Wang et al. (2023) under $\ell_{2}$ perturbations ($\epsilon=0.5$).
-Analysis on the effect of different parameter on robust accuracy
-Figure 4 illustrates how the robust accuracy of WDA varies as the parameter $\kappa$ increases.
-Across all models, raising $\kappa$ beyond 1 generally leads to a noticeable drop in robust accuracy.
-Specifically, for the models from Wang et al.
-(2023), Sehwag et al.
-(2022), and Sehwag et al.
-(2022) (RN-18), the best performance is observed at $\kappa=3$, whereas for Ding et al.
-(2020), the lowest robust accuracy occurs at $\kappa=2$.
-These results indicate that increasing $\kappa$ consistently weakens model robustness, with the precise $\kappa$ that produces the largest drop depending on the architecture.
-Figure 5 presents robust accuracy for the Wang et al.
-(2023) model with $r_{2}$ perturbations and with different step sizes for attack.
-The x-axis represents iterations, and the y-axis shows robust accuracy.
-In Figure 5, smaller step sizes (0.01 and 0.005) lead to higher robust accuracy (97%–99%), reflecting weaker attacks.
-The most effective attack occurs at step size 0.2, where the accuracy drops to around 88%.
-At larger step sizes (0.5 and 1.0), robust accuracy stabilizes at lower values despite initial drops, suggesting reduced attack effectiveness.
+Results on CIFAR-100 and ImageNet.
+The dominance of WDA++ generalizes to larger-scale datasets.
+On CIFAR-100, while point-wise WDA ($\kappa=1$) remains competitive against APGD variants, WDA++ ($\Omega_{1}$) proves most effective, yielding the lowest single-method robustness.
+This trend continues on ImageNet, where WDA++ ($\Omega_{1}$) outperforms all baselines, demonstrating that the full distributional threat model offers the strongest evaluation of model robustness across architectures.
+Finally, the ensemble A3-WDA consistently sets the state-of-the-art benchmark among ensemble methods.
+Runtime comparison.
+Figure 4(a) benchmarks the scalability of the WDA framework against PGD baselines on ImageNet, exhibiting a linear relationship between runtime and sample size across all methods.
+WDA variants consistently maintain comparable or lower computational costs than their PGD counterparts.
+Epsilon effect.
+Figure 4(b) demonstrates that WDA++ degrades performance faster than W-PGD as $\epsilon$ increases.
+This stems from W-PGD’s suboptimal use of global projection, which uniformly rescales perturbations and allows high-cost samples to dilute the budget.
+In contrast, WDA++ employs a searching method that prioritizes cheaper samples, saving budget from hard-to-attack inputs to facilitate larger, necessary perturbations on vulnerable ones.
+(a) Runtime comparison of different algorithms on the ImageNet dataset.
+(b) Robust accuracy of Sehwag et al. (2022)’s ResNet-18 on CIFAR-10 under increasing $\epsilon$.
 7 Conclusions
@@
 For ReLU networks, we derived exact WDRO bounds using their piecewise-affine property, computing data-dependent Lipschitz constants from activation patterns that significantly tighten existing global bounds.
-For networks with smooth activations (GELU, SiLU), we characterized the worst-case loss through asymptotic Jacobian behavior along adversarial rays, providing the first tractable WDRO analysis for these modern architectures.
-Our Wasserstein Distributional Attack (WDA) algorithm constructs adversarial distributions on $2N$ points rather than restricting to $N$ perturbed points, achieving lower robust accuracy than state-of-the-art methods across CIFAR-10/100 benchmarks.
-While WDA incurs additional computational overhead compared to single-point attacks due to evaluating multiple candidate perturbations per iteration, it demonstrates that existing robustness evaluations significantly underestimate vulnerability by considering only point-wise perturbations.
+For networks with smooth activations (e.g., GELU, SiLU), we characterized the worst-case loss through asymptotic Jacobian behavior along adversarial rays, providing the first tractable WDRO analysis for these modern architectures.
+Our Wasserstein Distributional Attacks (WDA, WDA++) construct adversarial distributions on $2N$ points rather than restricting to $N$ perturbed points, achieving lower robust accuracy than state-of-the-art methods across CIFAR-10/100 and ImageNet benchmarks.
+While WDA and WDA++ incur additional computational overhead compared to single-point attacks due to evaluating multiple candidate perturbations per iteration, it demonstrates that existing robustness evaluations significantly underestimate vulnerability by considering only point-wise perturbations.
 Together, these contributions narrow the gap between theoretical certificates and practical evaluation, revealing that both tighter bounds and stronger attacks emerge from properly leveraging network geometry and distributional perspectives.
-Appendix A Proofs
-A.1 Proof of Theorem 3.1
-Proof of Upper Bound
+Appendix A Preliminaries on WDRO
+Recall that given two probability measures $\mathbb{P}$ and $\mathbb{Q}$ on $\mathcal{Z}$, the Wasserstein distance is defined as
+$\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\left(\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{Z}\times\mathcal{Z}}d^{p}(z^{\prime},z)\,\mathrm{d}\pi(z^{\prime},z)\right)^{1/p}$
+for $p\in[1,\infty)$, and $\mathcal{W}_{d,p}(\mathbb{P},\mathbb{Q})\triangleq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\operatorname{ess.sup}_{{\pi}}(d)$ for $p=\infty$, where the feasible set is given by
+$\Pi(\mathbb{P},\mathbb{Q})\triangleq\bigl\{\pi\text{ on }\mathcal{Z}\times\mathcal{Z}\mid\pi(A\times\mathcal{Z})=\mathbb{P}(A),\;\pi(\mathcal{Z}\times B)=\mathbb{Q}(B)\ \forall A,B\subseteq\mathcal{Z}\bigr\}$
+the set of couplings (transport plans) between $\mathbb{P}$ and $\mathbb{Q}$.
+Intuitively, a transportation plan $\pi$ is feasible if it is a joint distribution whose first marginal is $\mathbb{P}$ and second marginal $\mathbb{Q}$.
+In the ambiguity set $\Omega_{p}=\left\{\mathbb{P}\mid\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\right\}$ of problem equation 3, a (distributional) attack $\mathbb{P}$ is admissible if the minimal effort for moving mass from $\mathbb{P}$ to the empirical distribution $\mathbb{P}_{N}$ is not exceeding budget $\epsilon$.
+Unlike traditional approaches which only allows point-wise perturbations, WDRO min-max model equation 3 allows both discrete and continuous distribution $\mathbb{P}$, which is extremely practical in certain scenarios where the ground-truth distribution $\mathbb{P}_{\rm true}$ is unknown and possibly continuous.
+Appendix B Proofs of Main Results
+B.1 Proof of Theorem 3.1
+Proof of Upper Bound.
 It is a standard result that for any $y=\bm{e}_{k}$, if $\ell$ is the cross-entropy loss then
-$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\log\left[\operatorname{softmax}\theta(x^{\prime})\right]_{k}-\log\left[\operatorname{softmax}\theta(x)\right]_{k}\right|\\ &\leq 2^{1/s}\|\theta(x^{\prime})-\theta(x)\|_{s},\end{array}$ (11)
+$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\log\left(\operatorname{softmax}\theta(x^{\prime})\right)_{k}-\log\left(\operatorname{softmax}\theta(x)\right)_{k}\right|\\ &\leq 2^{1/r}\|\theta(x^{\prime})-\theta(x)\|_{s},\end{array}$ (12)
 or if $\ell$ is the DLR loss then
-$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\left(\max_{k_{1}\neq k}\theta(x^{\prime})_{k_{1}}-\theta(x^{\prime})_{k}\right)-\left(\max_{k_{2}\neq k}\theta(x)_{k_{2}}-\theta(x)_{k}\right)\right|\\ &\leq 2^{1/s}\|\theta(x^{\prime})-\theta(x)\|_{s}.\end{array}$ (12)
-In addition, by Jordan & Dimakis (2020), we have that for any $x^{\prime},x\in\mathcal{X}$,
+$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&=\left|\left(\max_{k_{1}\neq k}\theta(x^{\prime})_{k_{1}}-\theta(x^{\prime})_{k}\right)-\left(\max_{k_{2}\neq k}\theta(x)_{k_{2}}-\theta(x)_{k}\right)\right|\\ &\leq 2^{1/r}\|\theta(x^{\prime})-\theta(x)\|_{s}.\end{array}$ (13)
+In addition, by Jordan and Dimakis (2020), we have that for any $x^{\prime},x\in\mathcal{X}$,
 $\|\theta(x^{\prime})-\theta(x)\|_{s}\leq\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}\times\|x^{\prime}-x\|_{r}.$ (14)
 Thus,
-$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&\leq 2^{1/s}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}\times\|x^{\prime}-x\|_{r}\\ &=\bm{L}\times d((x^{\prime},y),(x,y)).\end{array}$ (14)
-for any $x^{\prime},x\in\mathcal{X}$ and therefore by using Lipschitz certificate (Mohajerin Esfahani & Kuhn, 2018; Blanchet et al., 2019; Gao & Kleywegt, 2023; Gao et al., 2024; Chu et al., 2024), we have
+$\begin{array}[]{ll}|\ell(x^{\prime},y;\theta)-\ell(x,y;\theta)|&\leq 2^{1/r}\max_{\bm{D}\in\mathcal{D}_{\mathcal{X}}}\left\|J_{\bm{D}}\right\|_{{r}\to{s}}\times\|x^{\prime}-x\|_{r}\\ &=\bm{L}\times d((x^{\prime},y),(x,y)).\end{array}$ (15)
+for any $x^{\prime},x\in\mathcal{X}$ and therefore by using Lipschitz certificate (Mohajerin Esfahani and Kuhn, 2018; Blanchet et al., 2019; Gao and Kleywegt, 2023; Gao et al., 2024; Chu et al., 2024), we have
 $\sup_{\mathbb{P}\colon\mathcal{W}_{d,1}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon}\mathbb{E}_{\mathbb{P}}[\ell(Z;\theta)]\leq\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{L}\epsilon,$ (16)
 for any $\epsilon>0$.
-Proof of Lower Bound
+Proof of Lower Bound.
 To show that the lower bound of the worst-case loss is $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$, it is enough to construct a perturbation $\tilde{Z}$, a weight $\eta\in(0,1]$, and a distribution
@@
 so that $\mathcal{W}_{d,1}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$ and $\mathbb{E}_{\mathbb{P}_{\mathrm{adv}}}[\ell(Z;\theta)]\approx\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$.
-Since $\mathcal{D}_{\mathcal{X}}$ is finite, let $x^{\star},\bm{D}^{\star},k^{\prime\star},k^{\star}$ and sequence $\{u^{\star}_{t}\}$ be the maximizer in 7, i.e., $\bm{D}^{\star}\in\mathcal{D}_{x^{\star}},k^{\prime\star}\neq k^{\star},\{u^{\star}_{t}\}\subset\operatorname{int}(\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}}))$ and
+Since $\mathcal{D}_{\mathcal{X}}$ is finite, let $x^{\star},\bm{D}^{\star},k^{\prime\star},k^{\star}$ and sequence $\{u^{\star}_{t}\}$ be the maximizer in equation 7, i.e., $\bm{D}^{\star}\in\mathcal{D}_{x^{\star}},k^{\prime\star}\neq k^{\star},\{u^{\star}_{t}\}\subset\operatorname{int}(\operatorname{rec}(\mathcal{C}_{\bm{D}^{\star}}))$ and
 $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})^{\top}J_{\bm{D}^{\star}}u^{\star}_{t}\rightarrow\bm{l}\text{ when }t\rightarrow\infty.$
@@
 $\nabla_{x}\theta(\tilde{x})=J_{\bm{D^{\star}}},\quad\theta(\tilde{x})-\theta(x^{\star})=\alpha J_{\bm{D^{\star}}}u^{\star}_{t}.$ (19)
-Choose root $\iota$ so that $(X^{(\iota)},Y^{(\iota)}=\bm{e}_{k^{\star}})$. Then when $\ell$ is the cross-entropy loss or DLR loss, by a technical Lemma A.1 one has
+Choose root $\iota$ so that $(X^{(\iota)},Y^{(\iota)}=\bm{e}_{k^{\star}})$. Then when $\ell$ is the cross-entropy loss or DLR loss, by a technical Lemma C.1 one has
 $\lim_{\alpha\to\infty}\frac{\Delta\ell(\alpha)}{\alpha}=\lim_{\alpha\to\infty}\frac{\ell(\tilde{x},Y^{(\iota)};\theta)-\ell(x^{\star},Y^{(\iota)};\theta)}{\alpha}\geq v_{k^{\prime\star}}-v_{k^{\star}}.$ (20)
@@
 Therefore, the lower bound of the worst-case loss is $\mathbb{E}_{\mathbb{P}_{N}}[\ell(Z;\theta)]+\bm{l}\epsilon$.
-Sufficient condition of $\bm{l}=\bm{L}$
-Suppose that the dual-norm maximizer $\xi=\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of 7, then we have
+Figure 5: Illustration of Proof of Lower Bound
+Sufficient condition of $\bm{l}=\bm{L}$.
+Suppose that the dual-norm maximizer $\xi=\mathcal{M}_{r}(J_{\bm{D}^{\star}}^{\top}(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}}))\in\mathrm{rec}({\mathcal{C}_{\bm{D}^{\star}}})$ where $\bm{D}^{\star}$ is a maximizer of equation 6 and $(k^{\prime\star},k^{\star})$ is a maximizer of equation 7, then we have
 $\bm{l}$ $=(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})^{\top}J_{\bm{D}^{\star}}u^{\star}_{t}$ (24)
@@
 $=\|(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})\|_{r}\|J_{\bm{D}^{\star}}\|_{r\rightarrow s}$
-$=2^{1/s}\|J_{\bm{D}^{\star}}\|_{r\rightarrow s}=\bm{L},$
+$=2^{1/r}\|J_{\bm{D}^{\star}}\|_{r\rightarrow s}=\bm{L},$
 where the second last equality holds true because $(\bm{e}_{k^{\prime\star}}-\bm{e}_{k^{\star}})$ is the largest increment direction of $J_{\bm{D}^{\star}}$.
-Figure 6: Illustration of Proof of Lower Bound
-A.2 Proof of Theorem 3.3
+B.2 Proof of Theorem 3.3
 Proof.
 The proof for the upper and lower bounds is similar to the methodology we discussed in our previous exchange.
-Proof of Upper Bound The WDRO upper bound is a direct consequence of the Lipschitz continuity of the loss function.
+Proof of Upper Bound.
+The WDRO upper bound is a direct consequence of the Lipschitz continuity of the loss function.
 The Lipschitz constant of the combined loss function, $L_{\ell}=\sup_{Z\in\mathcal{Z}}\|\nabla_{x}\ell(Z;\theta)\|_{r}$, is bounded by the product of the Lipschitz constant of the loss with respect to the output and the Lipschitz constant of the network.
 That is,
-$\bm{L}\leq\|J(x)\|_{r\to s}\cdot\|\nabla_{\theta}\ell\|_{s}=\|J(x)\|_{r\to s}\cdot 2^{1/s}$
-Proof of Lower Bound The proof of the lower bound is identical with the ReLU network case, where it relies on constructing a specific adversarial distribution.
+$\bm{L}\leq\|J_{\theta}(x)\|_{r\to s}\cdot\|\nabla_{\theta}\ell\|_{s}=\|J_{\theta}(x)\|_{r\to s}\cdot 2^{1/r}$
+Proof of Lower Bound.
+The proof of the lower bound is identical with the ReLU network case, where it relies on constructing a specific adversarial distribution.
 This finds a point $x^{\star}$ and a direction $u^{\star}$ that maximize the rate of change of the loss.
@@
 ∎
-A.3 Technical Proofs
-Lemma A.1 (Technical lemma).
-In equation 19, if $\ell=\ell_{CE}$ is the cross-entropy loss, then
+B.3 Proof of Corollary 3.4
+Proof.
+Transformer blocks are compositions of differentiable primitives (linear maps, softmax attention, residual connections, LayerNorm with $\varepsilon>0$, and smooth activations), hence differentiable on $\mathcal{X}$.
+To obtain equation 10, one combines explicit Jacobian bounds for softmax self-attention from prior work (Kim et al., 2021; Castin et al., 2024) with standard Lipschitz calculus: for compositions, operator norms multiply (chain rule), and for residual maps $x\mapsto x+f(x)$ one has $\mathrm{Lip}(\mathrm{Id}+f)\leq 1+\mathrm{Lip}(f)$.
+Applying these rules block-by-block yields a product-form upper bound for $\sup_{x\in\mathcal{X}}|J_{\theta}(x)|_{2\to 2}$, and the general $r\to s$ case follows from Hölder’s inequality.
+∎
+Appendix C Auxiliary Results
+Lemma C.1 (Asymptotics of loss).
+If $\ell=\ell_{CE}$ is the cross-entropy loss in equation 20, then
 $\lim_{\alpha\to\infty}\frac{\Delta\ell_{CE}}{\alpha}=\max_{i}(J_{\bm{D}^{\star}}u^{\star}_{t})_{i}-(J_{\bm{D}^{\star}}u^{\star}_{t})_{k^{\star}}.$
@@
 We will analyze the limit for each loss function separately.
-Cross-Entropy Loss: The difference in loss is $\Delta\ell_{CE}=\ell_{CE}(\theta(\tilde{x}),e_{k^{\star}})-\ell_{CE}(\theta(x^{\star}),e_{k^{\star}})$.
+Cross-Entropy Loss.
+The difference in loss is $\Delta\ell_{CE}=\ell_{CE}(\theta(\tilde{x}),e_{k^{\star}})-\ell_{CE}(\theta(x^{\star}),e_{k^{\star}})$.
 Using the property $\ell_{CE}(z,e_{k^{\star}})=-(z_{k^{\star}}-\log\sum_{k}e^{z_{k}})$, the loss difference is:
@@
 $v_{\max}-v_{k^{\star}}=\max_{k}(J_{\bm{D}^{\star}}u^{\star}_{t})_{k}-(J_{\bm{D}^{\star}}u^{\star}_{t})_{k^{\star}}$
-DLR Loss: The difference in DLR loss is $\Delta\ell_{DLR}=\ell_{DLR}(\theta(\tilde{x}),k^{\star})-\ell_{DLR}(\theta(x^{\star}),k^{\star})$.
+DLR Loss.
+The difference in DLR loss is $\Delta\ell_{DLR}=\ell_{DLR}(\theta(\tilde{x}),k^{\star})-\ell_{DLR}(\theta(x^{\star}),k^{\star})$.
 $\Delta\ell_{DLR}=\left(\max_{k\neq k^{\star}}\theta(\tilde{x})_{k}-\theta(\tilde{x})_{k^{\star}}\right)-\left(\max_{k\neq k^{\star}}\theta(x^{\star})_{k}-\theta(x^{\star})_{k^{\star}}\right)$
@@
 $\lim_{\alpha\to\infty}\frac{\Delta\ell_{DLR}}{\alpha}=\left(\max_{k\neq k^{\star}}v_{k}\right)-v_{k^{\star}}=\max_{k\neq k^{\star}}(J_{\bm{D}^{\star}}u^{\star}_{t})_{k}-(J_{\bm{D}^{\star}}u^{\star}_{t})_{k^{\star}}$
-Appendix B Additional result
-Figure 7: Comparison of step size effect on model Cui et al. (2024) under $r_{\infty}$ perturbation ($\epsilon=8/255$).
-Figure 7 illustrates the convergence behavior of Cui et al.
-(2024) in terms of robust accuracy.
-Larger step sizes (0.1, 0.5) lead to higher final accuracy, whereas a smaller step size of 0.02 results in the lowest robust accuracy, indicating the most effective attack.
-Figure 8: Comparison of robust accuracy with varying $\kappa$
-Figure 8 presents the robust accuracy across different values of $\kappa$.
-Among the tested settings, $\kappa=2$ consistently produces the lowest robust accuracy for all models (Cui et al.
-(2024), Wang et al.
-(2023), and Sehwag et al.
-(2022)).
-Appendix C Usage of Large Language Models (LLMs)
-We used ChatGPT solely for revising the writing of the paper.
-Note that revision here strictly means enhancing the clarity and readability of the text (e.g., fixing typos or constructing latex tables), and not for any other purposes.
+Corollary C.2 (Feasibility).
+Let $\mathbb{P}_{N}=\sum_{i=1}^{N}\mu_{i}\bm{\delta}_{(x^{(i)},y^{(i)})}$ and $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$.
+For $\alpha_{i}\in[0,1]$ and points $\tilde{x}^{(i)}\in\mathcal{X}$, define
+$\mathbb{P}_{\rm adv}\triangleq\sum_{i=1}^{N}\mu_{i}\Bigl((1-\alpha_{i})\bm{\delta}_{(x^{(i)},y^{(i)})}+\alpha_{i}\bm{\delta}_{(\tilde{x}^{(i)},y^{(i)})}\Bigr).$
+Then for any $p\in[1,\infty)$,
+$\mathcal{W}_{d,p}(\mathbb{P}_{\rm adv},\mathbb{P}_{N})^{p}\leq\sum_{i=1}^{N}\mu_{i}\alpha_{i}\|\tilde{x}^{(i)}-x^{(i)}\|_{r}^{p}.$
+In particular, if $\sum_{i=1}^{N}\mu_{i}\alpha_{i}\|\tilde{x}^{(i)}-x^{(i)}\|_{r}^{p}\leq\epsilon^{p}$, then $\mathbb{P}_{\rm adv}\in\Omega_{p}(\epsilon)=\{\mathbb{P}:\mathcal{W}_{d,p}(\mathbb{P},\mathbb{P}_{N})\leq\epsilon\}$.
+Proof.
+Use the explicit coupling that leaves $(1-\alpha_{i})$ mass of each atom fixed and transports $\alpha_{i}$ mass from $(\tilde{x}^{(i)},y^{(i)})$ to $(x^{(i)},y^{(i)})$. ∎
+Appendix D WDA Study
+The $\kappa$ parameter.
+Figure 6 illustrates how the robust accuracy of WDA varies as the parameter $\kappa$ increases.
+Across all models, raising $\kappa$ beyond 1 generally leads to a noticeable drop in robust accuracy.
+Specifically, for the models from Figure 6(a) in $r_{2}$, $\epsilon=0.5$ settings, the best performance is observed at $\kappa=3$, whereas for Ding et al.
+(2020); Cui et al.
+(2024); Wang et al.
+(2023); Sehwag et al.
+(2022) in Figure 6(b), the lowest robust accuracy occurs at $\kappa=2$.
+These results indicate that increasing $\kappa$ can weaken model robustness, with the precise $\kappa$ that produces the largest drop depending on the architecture.
+(a) CIFAR-10 – $r_{2}$, $\epsilon=0.5$.
+(b) CIFAR-10 – $r_{\infty}$, $\epsilon=8/255$.
+Figure 6: Analysis of robust accuracy across different $\kappa$ scales.
+The step size parameter.
+In Figure 7(a) ($r_{2},\epsilon=0.5$), we observe that step sizes significantly larger than the budget (e.g., 1.0 and 0.5) cause the fastest initial drop in accuracy but risk overshooting the optimal perturbation.
+In contrast, a moderate step size of 0.2 provides the most stable convergence to the lowest robust accuracy.
+Very small step sizes (0.01, 0.005) fail to explore the $\epsilon$-ball effectively within the given iterations, resulting in an underestimation of the attack’s potency.
+Similarly, in Figure 7(b) ($r_{\infty},\epsilon=8/255\approx 0.031$), the choice of step size is critical for convergence.
+A step size of 0.02 (approximately $2/3\epsilon$) yields the most potent attack.
+Step sizes significantly larger than $\epsilon$ (0.1, 0.5) lead to higher final robust accuracy due to oscillation around the boundary of the $L_{\infty}$ ball, while the smallest step size (0.005) suffers from insufficient progress, failing to reach the floor within 20 iterations.
+(a) CIFAR-10 – $r_{2}$, $\epsilon=0.5$.
+(b) CIFAR-10 – $r_{\infty}$, $\epsilon=8/255$.
+Figure 7: Analysis of convergence and step size effects across different perturbation types. (a) Method from Wang et al. (2023); (b) Method from Cui et al. (2024).
+Appendix E WDA++ Study
+While the standard WDA fixes a global scalar $\kappa$ for all samples, WDA++ adopts an adaptive strategy to allocate the transport budget more efficiently.
+The core intuition is to assign higher transport mass (larger $1/\kappa_{i}$) to samples that are geometrically closer to the decision boundary, as these require less “cost” to flip.
+We report the classification accuracy on the constructed distribution $\mathbb{P}_{\mathrm{adv}}$ as
+$\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa_{i}}\right)\times\bm{1}_{\{h_{\theta}(x_{i})=y_{i}\}}+\frac{1}{\kappa_{i}}\times\bm{1}_{\{h_{\theta}(x^{\mathrm{adv}}_{i})=y_{i}\}},$
+where $h_{\theta}(x)=\arg\max_{k}\theta_{k}(x)$ is the class predictor.
+The practical implementation involves three specific algorithmic components to ensure efficiency and optimality under the budget constraint.
+These are detailed below and correspond to the steps in Algorithm 2.
+Algorithm 2 WDA++
+Inputs: neural network $\theta:\mathbb{R}^{n}\to\mathbb{R}^{K}$, empirical distribution $\mathbb{P}_{N}=\frac{1}{N}\sum_{i=1}^{N}\bm{\delta}_{(X^{(i)},Y^{(i)})}$, budget $\epsilon>0$, cost-norm $r\in\{1,2,\infty\}$, Wasserstein order $p\in\{1,2\}$, step size $\alpha>0$, srchiter, maxiter, and topK
+Outputs: adversary distribution $\mathbb{P}_{\mathrm{adv}}$ such that $\mathcal{W}_{d,p}(\mathbb{P}_{\mathrm{adv}},\mathbb{P}_{N})\leq\epsilon$, where $d((x^{\prime},y^{\prime}),(x,y))=\|x^{\prime}-x\|_{r}+\infty\cdot\bm{1}_{\{y^{\prime}\neq y\}}$
+Initialize: dual-norm maximizer $\mathcal{M}$ equation 1
+for $i=1$ to NN do
+  if $\arg\max_{c\in\{1,\dots,K\}}\theta_{c}(x_{0})\neq k$ then $X_{\mathrm{adv}}^{(i)}\leftarrow X^{(i)}$, $d_{i}\leftarrow 0$ continue
+  $x_{0}\leftarrow X^{(i)}$
+  $\bm{e}_{k}\leftarrow Y^{(i)}$ for some $k\in\{1,\dots,K\}$
+  $\mathcal{J}_{i}\leftarrow\arg\max_{\begin{subarray}{c}\mathcal{J}\subseteq\{1,\dots,K\}\setminus\{k\}\\ |\mathcal{J}|=\texttt{topK}\end{subarray}}\sum_{j\in\mathcal{J}}\theta_{j}(x_{0})$
+  for iter =0=0 to maxiter do
+   $g_{j}\leftarrow\nabla_{x}\theta(x_{\mathrm{iter}})^{\top}(\bm{e}_{j}-\bm{e}_{k})$ for $j\in\mathcal{J}_{i}$
+   $u_{j}\leftarrow\mathcal{M}_{r}(g_{j})$ for $j\in\mathcal{J}_{i}$
+   $\varphi_{j}\leftarrow x_{\mathrm{iter}}+\alpha u_{j}$ for $j\in\mathcal{J}_{i}$
+   $j^{*}\leftarrow\arg\max_{j\in\mathcal{J}_{i}}\left(\theta_{j}(\varphi_{j})-\theta_{k}(\varphi_{j})\right)$
+   $x_{\mathrm{iter+1}}\leftarrow\varphi_{j^{*}}$
+   if $\arg\max_{c\in\{1,\dots,K\}}\theta_{c}(x_{\mathrm{iter+1}})\neq k$ then $x_{\mathrm{pre}}^{(i)}\leftarrow x_{\mathrm{iter}}$, $x_{\mathrm{flip}}^{(i)}\leftarrow x_{\mathrm{iter+1}}$ break
+  end for
+  if $\arg\max_{c\in\{1,\dots,K\}}\theta_{c}(x_{\mathrm{iter+1}})=k$ then $X_{\mathrm{adv}}^{(i)}\leftarrow X^{(i)}$, $d_{i}\leftarrow+\infty$ continue
+  $\Delta\leftarrow x_{\mathrm{flip}}^{(i)}-x_{\mathrm{pre}}^{(i)}$, $t_{\ell}\leftarrow 0$, $t_{u}\leftarrow 1$
+  for $j=1$ to srchiter do
+   $t\leftarrow(t_{\ell}+t_{u})/2$
+   $\tilde{x}\leftarrow x_{\mathrm{pre}}^{(i)}+t\Delta$
+   if $\arg\max_{c}\theta_{c}(\tilde{x})\neq k$ then $t_{u}\leftarrow t$ else $t_{\ell}\leftarrow t$
+  end for
+  $X_{\mathrm{adv}}^{(i)}\leftarrow x_{\mathrm{pre}}^{(i)}+t_{u}\Delta$
+  $d_{i}\leftarrow\|X_{\mathrm{adv}}^{(i)}-X^{(i)}\|_{r}$
+end for
+$\mathcal{I}\leftarrow\arg\mathrm{sort}\left(\left\{d_{i}:i\in\{1,\dots,N\}\right\}\right)$
+$\kappa_{i}\leftarrow+\infty$ for $i\in\{1,\dots,N\}$
+$B\leftarrow\epsilon^{p}$
+for $i\in\mathcal{I}$ do
+  if $B\leq 0$ then break
+  $\kappa_{i}\leftarrow{1}/{\min\left\{1,\ {NB}/{d_{i}^{p}}\right\}}$
+  $B\leftarrow B-\frac{1}{N}\frac{1}{\kappa_{i}}d_{i}^{p}$
+end for
+$\mathbb{P}_{\mathrm{adv}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{1}{\kappa_{i}}\right)\bm{\delta}_{(X^{(i)},Y^{(i)})}+\frac{1}{\kappa_{i}}\bm{\delta}_{(X_{\mathrm{adv}}^{(i)},Y^{(i)})}$
+return $\mathbb{P}_{\mathrm{adv}}$
+Direction search via Top-K heuristic.
+Evaluating the gradient for all K1K-1 target classes can be computationally expensive for datasets with many classes (e.g., CIFAR-100, ImageNet).
+To accelerate the attack, we employ a topK heuristic.
+We restrict the search for the optimal adversarial direction to the top-$K$ classes with the highest logits (excluding the correct class).
+In our experiments, we set $\texttt{topK}=20$ for ImageNet, $\texttt{topK}=10$ for CIFAR-100 and $\texttt{topK}=5$ for CIFAR-10.
+Boundary estimation via binary search.
+To compute the minimal flip cost, we cannot rely solely on the gradient step, as it may overshoot or undershoot the decision boundary.
+Instead, we perform a binary search along the direction of the successful perturbation vector $\Delta=x_{\mathrm{flip}}^{(i)}-x_{\mathrm{pre}}^{(i)}$.
+This refinement step, controlled by $\texttt{srchiter}=10$, gives $d_{i}$.
+Budget allocation.
+WDA++ adaptively allocates the global Wasserstein budget $\epsilon$ by treating the attack as a Knapsack-like optimization problem: maximizing misclassified samples subject to $\sum\mu_{i}d_{i}^{p}/\kappa_{i}\leq\epsilon^{p}$.
+We adopt an optimal greedy strategy that calculates the flip cost $d_{i}$ for each sample, sorts them in ascending order, and sequentially assigns maximum weight (up to $\kappa_{i}=1$) to the “cheapest” samples until the budget $B=\epsilon^{p}$ is exhausted.
+This prioritizes the weakest links in the model to maximize the error rate for a given distributional radius.